### v1

# Adaptive and oblivious statistical adversaries are equivalent

Abstract
We resolve a fundamental question about the ability to perform a statistical task, such as learning, when an adversary corrupts the sample. Such adversaries are specified by the types of corruption they can make and their level of knowledge about the sample. The latter distinguishes between sample-adaptive adversaries which know the contents of the sample when choosing the corruption, and sample-oblivious adversaries, which do not. We prove that for all types of corruptions, sample-adaptive and sample-oblivious adversaries are equivalent up to polynomial factors in the sample size. This resolves the main open question introduced by [BLMT22] and further explored in [CHL+23].
Specifically, consider any algorithm $A$ that solves a statistical task even when a sample-oblivious adversary corrupts its input. We show that there is an algorithm $A^{\prime}$ that solves the same task when the corresponding sample-adaptive adversary corrupts its input. The construction of $A^{\prime}$ is simple and maintains the computational efficiency of $A$: It requests a polynomially larger sample than $A$ uses and then runs $A$ on a uniformly random subsample.
One of our main technical tools is a new structural result relating two distributions defined on sunflowers which may be of independent interest.
1 Introduction
Classic models of data analysis assume that data is drawn independently from the distribution of interest, but the real world is rarely so kind. To be robust to the messiness of real-world data, we desire algorithms that succeed even in the presence of an adversary that corrupts the data. Such adversaries were first introduced in the seminal works of [Tuk60, Hub64, Ham71] and have since been the subject of intense study in a variety of settings [Val85, Hau92, KL93, KSS94, BEK02, DSFT+14, LRV16, CSV17, DKK+19, DKPZ21, HSSVG22, BLMT22, CHL+23, DK23].
By now, there are numerous models for how the adversary can corrupt the data, including additive, subtractive, “strong”/“nasty”, agnostic, and adaptive and non-adaptive variants of each of these. In many cases, the provable guarantees for our algorithms are only known for a subset of these models. We refer the interested reader to the excellent recent textbook [DK23] for a more complete background and survey of recent results and open directions. Our work focuses on a surprisingly under-explored question:
What is the relationship between the various statistical adversaries?
Specifically, we compare adaptive adversaries, which can look at the sample before deciding on a corruption, and oblivious adversaries, which must commit to their corruptions before the i.i.d. sample is drawn.
Theorem 1.
Adaptive adversaries and their oblivious counterparts are equivalent up to scaling the sample size by a factor polynomial in the original sample size and polylogarithmic in the domain size.
Theorem 1 resolves the main question introduced by [BLMT22] and further explored in [CHL+23]. We defer its formal statement to Section 2, but, for now, mention two points. First, it is a generic result that proves the equivalency between many distinct adaptive adversaries and their oblivious counterparts (e.g. the equivalence between “subtractive adaptive” and “subtractive oblivious” adversaries). Second, it is constructive. We give a simple transformation, the subsampling filter described in Definition 5, which takes any algorithm that succeeds on a statistical task in the presence of the oblivious adversary and converts it to one that succeeds on the same task in the presence of the adaptive adversary. This transformation preserves the statistical and computational efficiency of the original algorithm up to polynomial factors.
In addition to answering a foundational question about the relative power of statistical adversaries, Theorem 1 has several practical implications:
  1. 1.
    Given the many distinct definitions of robustness, it can be difficult for a practitioner to determine which definition is most appropriate for their setting and therefore which algorithm to utilize. Theorem 1 partially alleviates this issue by greatly reducing the number of truly unique adversary models.
  2. 2.
    It shows that a single algorithmic idea, that of subsampling, amplifies robustness in many different models. Formally, it takes an algorithm that is only robust to the oblivious adversary and converts it to one robust to the adaptive counterpart. This suggests that, even if the practitioner cannot precisely determine the most appropriate model of robustness, they should try subsampling.
  3. 3.
    Theorem 1 can be reformulated as an answer to an equivalent and independently interesting question: How useful is it to hide one’s dataset from the adversary? It shows that private data does not afford much more robustness than public data.
2 Our Results
Before formally describing our main result, we define a unified framework in which to express and analyze statistical adversaries. It may be instructive to view this framework with the following concrete adaptive adversary and its oblivious counterpart in mind:
Example 1.
Consider the following adaptive and oblivious adversaries parameterized by $\eta\in[0,1]$:
  • Adaptive: When the algorithm requests $n$ points, first an i.i.d. sample $\bm{S}\sim\mathcal{D}^{n}$ is drawn. Then, the adversary chooses $\lfloor\eta\cdot n\rfloor$ arbitrary points in the sample that it can remove and replace with arbitrarily chosen points. The algorithm receives this corrupted sample.
  • Oblivious: The adversary can choose any $\mathcal{D}^{\prime}$ that has a total variation distance to $\mathcal{D}$ of at most $\eta$, and the algorithm receives $n$ i.i.d. draws from $\mathcal{D}^{\prime}$.
These two adversaries are well-studied, and are referred to by different names. The adaptive adversary is typically referred to as “strong contamination” in the statistical estimation literature [DK23] and “nasty noise” in the PAC learning literature [BEK02]. The oblivious adversary has been referred to as “general, non-adaptive, contamination”[DK23]. In our unified framework, these adversaries will be defined via the same “cost-function,” and as a result, we prove them equivalent.
2.1 A unified framework to define statistical adversaries
Each adversary will be parameterized by a “cost” function $\rho$ where $\rho(x,y)$ specifies the cost the adversary pays to corrupt $x$ to $y$, with a cost of $\infty$ indicating that the adversary is not allowed to change $x$ to $y$. The adversary can choose any corruptions subject to a budget constraint on the total cost incurred. This cost function is required to have two basic properties.
Definition 1 (Cost function).
A function $\rho:X\times X\to\mathds{R}_{\geq 0}\cup\{\infty\}$ is said to be a “cost function” if it satisfies the following properties.
  1. 1.
    For any $x\in X$, $\rho(x,x)=0$.
  2. 2.
    For any $x,y\in X$, $\rho(x,y)\geq 0$.
The adversary is specified by both the cost function and whether it is adaptive or oblivious. Given the cost function, $\rho$, the corresponding adaptive adversary is defined as follows:
Definition 2 (Adaptive adversary, corruptions to the sample).
For any cost function $\rho$ and $S\in X^{n}$, we use $\mathcal{C}_{\rho}(S)$ to denote all $S^{\prime}\in X^{n}$ for which
$\frac{1}{n}\sum_{i\in[n]}\rho(S_{i},S^{\prime}_{i})\leq 1.$
The $\rho$-adaptive adversary is allowed to corrupt the clean sample $S$ to any $S^{\prime}\in\mathcal{C}_{\rho}(S)$. For any $f:X^{n}\to\{0,1\}$ and distribution $\mathcal{D}$, the max success probability of $f$ in the presence of the $\rho$-adaptive adversary is denoted:
$\mathrm{Adaptive\text{-}Max}_{\rho}(f,\mathcal{D})\coloneqq\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{n}}\left[\sup_{\bm{S}^{\prime}\in\mathcal{C}_{\rho}(\bm{S})}\left\{f(\bm{S}^{\prime})\right\}\right].$
In the case of the adversaries of Example 1, their cost functions are simply $\rho(x,y)=1/\eta$ for all $x\neq y$. In that case, the budget constraint in the above definition ensures that, for this choice of cost function, $\rho$, the $\rho$-adaptive adversary can corrupt at most an $\eta$ fraction of points in the sample, corresponding to the standard definition of the “strong contamination”/“nasty noise” models.
Given a cost function, the associated oblivious adversary replaces the budget constraint of the adaptive setting with a natural distributional analog. It is easy to see that the following definition of $\rho$-oblivious adversaries is equivalent to the “general, non-adaptive, contamination” model of Example 1 when the cost function is defined as $\rho(x,y)=1/\eta$ for all $x\neq y$.
Definition 3 (Oblivious adversary, corruptions to a distribution).
For any cost function $\rho$ and distribution $\mathcal{D}$, we overload $\mathcal{C}_{\rho}(\mathcal{D})$ to refer to the set of all distributions $\mathcal{D}^{\prime}$ for which there exists a coupling of $\bm{x}\sim\mathcal{D}$ and $\bm{x}^{\prime}\sim\mathcal{D}^{\prime}$ satisfying
$\mathop{{\mathds{E}}\/}[\rho(\bm{x},\bm{x}^{\prime})]\leq 1.$
The $\rho$-oblivious adversary is allowed to corrupt the base distribution $\mathcal{D}$ to any $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$. For any $f:X^{n}\to\{0,1\}$ and distribution $\mathcal{D}$, the max success probability of $f$ in the presence of the $\rho$-oblivious adversary is denoted:
$\mathrm{Oblivious\text{-}Max}_{\rho}(f,\mathcal{D})\coloneqq\sup_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{\mathop{{\mathds{E}}\/}_{\bm{S}^{\prime}\sim(\mathcal{D}^{\prime})^{n}}\left[f(\bm{S}^{\prime})\right]\right\}.$
In Definition 3, since $\bm{x}\sim\mathcal{D}$ and $\bm{x}^{\prime}\sim\mathcal{D}^{\prime}$ can be coupled so that the average cost to corrupt $\bm{x}$ to $\bm{x}^{\prime}$ is at most $1$, we can similarly couple $\bm{S}\sim\mathcal{D}^{n}$ and $\bm{S}^{\prime}\sim(\mathcal{D}^{\prime})^{n}$ so that the average cost to corrupt each point in $\bm{S}$ to the corresponding point in $\bm{S}^{\prime}$ is at most $1$. From this perspective, the crucial difference between the oblivious and adaptive adversary is that the oblivious adversary must commit to how it corrupts each $\bm{x}$ without knowing the contents of the sample, whereas the adaptive adversary gets to view $\bm{S}$ before deciding.
We show, in Section 3, that our framework can express many commonly studied statistical adversaries, including subtractive contamination, additive contamination, and agnostic noise.
Remark 1 (Partially-adaptive statistical adversaries).
Some statistical adversaries lie between their fully adaptive and fully oblivious counterparts. These include malicious noise [Val85] and the non-iid oblivious adversary defined in [CHL+23]. Our results readily extend to such adversaries (see Section 3.1 for details).
2.2 Our main result: Adaptive and oblivious adversaries are equivalent
Our main result is that for any algorithm $A$ and cost function $\rho$, there exists an algorithm $A^{\prime}$ inheriting the efficiency of $A$ for which the performance of $A$ in the presence of the oblivious adversary is equivalent to the performance of $A^{\prime}$ in the presence of the adaptive adversary.
Definition 4 ($\varepsilon$-equivalent algorithms).
For any algorithms $A:X^{n}\to Y$ and $A^{\prime}:X^{m}\to Y$, we say that $A$ in the presence of the $\rho$-oblivious adversary is $\varepsilon$-equivalent to $A^{\prime}$ in the presence of the $\rho$-adaptive adversary if for any test function $T:Y\to\{0,1\}$ and distribution $\mathcal{D}$ supported on $X$,
$\left|\mathrm{Oblivious\text{-}Max}_{\rho}(T\circ A,\mathcal{D})-\mathrm{Adaptive\text{-}Max}_{\rho}(T\circ A^{\prime},\mathcal{D})\right|\leq\varepsilon.$
Colloquially, $A$ and $A^{\prime}$ are $\varepsilon$-equivalent if no test can distinguish their outputs with more than $\varepsilon$ probability. Note that while the above definition is about the maximum acceptance probability of $T$, it also applies to the test $\overline{T}\coloneqq 1-T$ and therefore the minimum acceptance probability of $T$ also must be approximately the same for $A$ and $A^{\prime}$.
The algorithm $A^{\prime}$ will run $A$ on a uniformly random subsample of its input.
Definition 5 (Subsampling filter).
For any $m\geq n$ we define the subsampling filter $\Phi_{m\to n}:X^{m}\to X^{n}$ as the (randomized) algorithm that given $S\in X^{m}$, returns a sample of $n$ points drawn uniformly without replacement from $S$.
Theorem 2 (Subsampling neutralizes the adaptivity in statistical adversaries).
For any algorithm $A:X^{n}\to Y$, $\varepsilon>0$, and cost function $\rho$, let $m=\mathrm{poly}(n,\ln|X|,1/\varepsilon)$ and $A^{\prime}\coloneqq A\circ\Phi_{m\to n}$. Then, $A$ in the presence of the $\rho$-oblivious adversary is $\varepsilon$-equivalent to $A^{\prime}$ in the presence of the $\rho$-adaptive adversary.
For constant $\varepsilon$, Theorem 2 says that if there is an algorithm $A$ solving a statistical task with an oblivious adversary taking as input $n\cdot\log|X|$ bits, there is an algorithm $A^{\prime}$ solving the same task with an adaptive adversary taking only polynomially more bits as input. Furthermore, if $A$ is computationally efficient, then $A^{\prime}$ is too.
Remark 2 (Continuous domains).
In many statistical problems, the domain is $\mathds{R}^{d}$. To apply Theorem 2 to an algorithm $A$ over continuous domains, we first discretize that domain to some $X\coloneqq\mathrm{disc}(\mathds{R})^{d}$ where the discretization depends on $A$. If $A$ requires $b$ bits of precision in each dimension, then $\log_{2}|X|=bd$, which is typically polynomial in $n$. For example, under the mild assumption that $A$ accesses the bits of each dimension sequentially, both $b$ and $d$ are upper bounded by the time complexity of $A$. In this setting, if the time complexity of $A$ is polynomial in $n$, then so is $\ln|X|$.
Theorem 2 is a special case of our main theorem in which the $|X|$ is replaced with the degree of the cost function, a measure of the number of corruptions the adversary can make for each input.
Definition 6 (Degree of a cost function).
For any cost function $\rho:X\times X\to\mathds{R}_{\geq 0}\cup\{\infty\}$, the degree of $\rho$ is defined as
$\deg(\rho)\coloneqq\sup_{x\in X}\left\{\text{The number of distinct }y\in X\text{ for which }\rho(x,y)\neq\infty\right\}.$
Theorem 3 (Main result, generalization of Theorem 2).
For any algorithm $A:X^{n}\to Y$, $\varepsilon>0$, and cost function $\rho$ with degree $d\geq 2$, let $m=O\left(\frac{n^{3}(\ln d)^{2}}{\varepsilon^{4}}\right)$ and $A^{\prime}\coloneqq A\circ\Phi_{m\to n}$. Then, $A$ in the presence of the $\rho$-oblivious adversary is $\varepsilon$-equivalent to $A^{\prime}$ in the presence of the $\rho$-adaptive adversary.
The degree is constant for many natural cost functions, such as the cost function corresponding to subtractive contamination. In these cases, Theorem 3 has no dependence on the domain size.
2.3 Lower bounds
Theorem 3 requires a polynomial increase of the sample size of $A^{\prime}$ relative to $A$. It is natural to wonder whether such an increase is necessary. The main results of [CHL+23] show it is.
Fact 2.1 ([CHL+23]).
For any $n\in\mathds{N}$, the task of Gaussian mean testing with appropriate parameters (depending on $n$) can be solved using $n$ samples in the presence of the oblivious additive adversary, but requires $\tilde{\Omega}(n^{4/3})$ in the presence of the adaptive additive adversary.
In the setting of Theorem 3, one difficulty in interpreting Fact 2.1 is that the cost function corresponding to the additive adversary has a large degree [^1] Since the domain for the task is over the continuous domain of $\mathds{R}^{d}$, technically this cost function has infinite degree. However, as we discussed in Remark 2, it makes more sense to think of the degree as $\approx 2^{d}$ in this setting, which happens to be exponential in the $n$ of Fact 2.1., and so it’s unclear if the increased sample size is to due a dependence on the degree or an innate required polynomial increase.
For example, the subtractive adversary (formally defined in Section 3) has a degree of $2$ because, for each point in the sample, it chooses between keeping that point or removing it. For this adversary, is a polynomial increase in sample size necessary? Our first lower bound gives a straightforward proof this is the case, even for a simple task.
Theorem 4 (A polynomial increase in sample size is necessary).
Let $\mathcal{D}$ be a distribution on $X=[m]\coloneqq\{1,\ldots,m\}$ that is promised to be uniform on some $X^{\prime}\subseteq X$. Then,
  1. 1.
    There is an algorithm that estimates $|X^{\prime}|$ using $n\coloneqq\tilde{O}(\sqrt{m})$ samples even with oblivious subtractive contamination.
  2. 2.
    Any algorithm that estimates $|X^{\prime}|$ to the same accuracy with adaptive subtractive contamination requires $\tilde{\Omega}(m)$ samples.
Theorem 4 implies that in the statement of Theorem 2, we must take $m$ polynomial larger than $n$. Next, we show that this $m$ must also depend polylogarthmically on a degree-like characteristic of the cost function.
Definition 7 (Budget-bounded degree).
For any cost function $\rho:X\times X\to\mathds{R}_{\geq 0}\cup\{\infty\}$ and $b\in\mathds{R}_{\geq 0}$, the $b$-bounded degree of $\rho$ is defined as
$\deg_{b}(\rho)\coloneqq\sup_{x\in X}\left\{\text{The number of distinct }y\in X\text{ for which }\rho(x,y)\leq b\right\}.$
This lower bound will make one assumption on $\rho$: that $\rho(x,y)\geq 1+\delta$ whenever $x\neq y$ for a small constant $\delta$. This corresponds to the adversary having a budget on how many points they can change and is satisfied by all of the well-studied models discussed in Section 3.
Theorem 5 (Dependence on $\ln\deg_{b}(\rho)$ is necessary).
For any constants $b,\delta>0$, large enough $n\in\mathds{N}$, and cost function $\rho$ for which $\rho(x,y)\geq 1+\delta$ whenever $x\neq y$, there is an algorithm $A:X^{n}\to\{0,1\}$ for which the following holds. If $A$ in the presence of the $\rho$-oblivious adversary is $(\varepsilon=0.9)$-equivalent to $A^{\prime}\coloneqq A\circ\Phi_{{m}\to n}$ in the presence of the $\rho$-adaptive adversary, then
$m\geq\tilde{\Omega}_{b,\delta}\left(n\cdot\ln\deg_{b}(\rho)\right).$
Comparing Theorems 3 and 5, for “reasonable” cost functions in which $\deg(\rho)\approx\deg_{1000}(\rho)$ and $\rho(x,y)\geq 1.001$ for all $x\neq y$, a domain-size independent result is possible precisely when the degree does not grow with $|X|$.
2.4 Relation to recent work
Recent work of Blanc, Lange, Malik, and Tan initiated a formal study of the relationship between adaptive adversaries and their oblivious counterparts [BLMT22]. They conjectured the equivalence of adaptive and oblivious statistical adversaries but only proved it in two special cases.
  1. 1.
    They showed that additive oblivious and additive adaptive adversaries are equivalent. Our result, which applies to all statistical adversaries, requires an entirely different approach. This is because, in some sense, the adaptive additive adversary is less adaptive than other adaptive adversaries. We elaborate on this point in Section C.1 and explain why [BLMT22]’s approach does not generalize to all adversaries.
  2. 2.
    They also showed that if a statistical query (SQ) algorithm is robust to an oblivious adversary, it can be upgraded to be robust to the corresponding adaptive adversary. The restriction to SQ algorithms greatly facilitates [BLMT22]’s analysis because we have a much better understanding of SQ algorithms than general algorithms. For example, the quality of the best SQ algorithm for a given task is captured by simple combinatorial measures [BFJ+94, Fel17]. In Section C.2, we further describe [BLMT22]’s SQ result and give advantages of our result even for algorithms that can be cast in the SQ framework.
Other recent work of Canonne, Hopkins, Li, Liu, and Narayanan tackled the equivalency of statistical adversaries from the other direction [CHL+23]. While we aim to show that distinct statistical adversaries are equivalent, they showed a separation: For the well-studied problem of Gaussian mean testing, an adaptive adversary requires polynomial more samples than the corresponding oblivious adversary (see Fact 2.1).
3 Instantiating common adversaries in our framework
Here, we show how to express many common statistical adversaries within our framework. For completeness, we include the “strong contamination/nasty noise” adversary of Example 1.
Strong contamination/nasty noise: As mentioned in Example 1, both the “strong contamination/nasty noise” adversary, that can arbitrarily replace an $\eta$ fraction of an i.i.d. sample, and the “general, non-adaptive, contamination” adversary, that can perturb the underlying distribution from which the sample is drawn by at most $\eta$ in total variation distance, correspond to adaptive and oblivious adversaries with the following cost function:
$\rho_{\mathrm{strong}}(x,y)\coloneqq\begin{cases}0&\text{if }x=y\\ \tfrac{1}{\eta}&\text{otherwise.}\end{cases}$
Agnostic learning [Hau92, KSS94]: Agnostic noise is a well-studied adversary [KKMS08, KK09, Fel10, DSFT+14, DKPZ21] specific to supervised learning problems, where each point in the sample is a pair $(x,y)$ of the input and its label. This adversary is allowed to change $\eta$ fraction of the labels but must keep the inputs unchanged. It corresponds to the cost function,
$\rho_{\mathrm{agnostic}}((x_{1},y_{1}),(x_{2},y_{2}))\coloneqq\begin{cases}\infty&\text{if }x_{1}\neq x_{2}\\ \tfrac{1}{\eta}&\text{if }x_{1}=x_{2}\text{ and }y_{1}\neq y_{2}\\ 0&\text{if }x_{1}=x_{2}\text{ and }y_{1}=y_{2}.\\ \end{cases}$
Agnostic learning typically refers to the $\rho$-oblivious adversary. It can be thought of as the learner receiving an i.i.d. sample of points of the form $(\bm{x},g(\bm{x}))$ where $g$ is close to the original target $f$ in the sense that
$\mathop{{\operatorname{{Pr}}}\/}_{\bm{x}}[f(\bm{x})\neq g(\bm{x})]\leq\eta.$
In the adaptive variant, first, a sample is drawn that is labeled by the true target function. Then, an adversary may corrupt $\eta$-fraction of the labels arbitrarily. This variant is sometimes referred to as nasty classification noise [BEK02].
Note that the cost function $\rho_{\mathrm{agnostic}}$ only has degree $2$ in binary classification settings. Theorem 3 therefore shows the equivalence between nasty classification noise and agnostic noise with no dependence on the domain size.
Subtractive contamination: In the adaptive variant of subtractive contamination, the adversary is allowed to remove $\lfloor\eta n\rfloor$ points from a size-$n$ sample. The algorithm receives the remaining $n-\lfloor\eta n\rfloor$ points.
In the oblivious variant [DK23], the algorithm receives i.i.d. samples from the distribution $\mathcal{D}$ conditioned on some event $E$ that occurs with probability $1-\eta$. This can be thought of as the adversary removing $\eta$-fraction of the distribution corresponding to when the event $E$ does not occur.
To fit subtractive contamination into our framework, we will augment the domain with a special element $\varnothing$, to indicate the adversary has removed this point. For the augmented domain $X^{\prime}\coloneqq X\cup\{\varnothing\}$, it uses the cost function,
$\rho_{\mathrm{sub}}(x,y)\coloneqq\begin{cases}0&\text{if }x=y\\ \frac{1}{\eta}&\text{if }x\neq y\text{ and }y=\varnothing\\ \infty&\text{otherwise}.\end{cases}$
Note that once again, this cost function has degree only $2$, so by Theorem 3, the $\rho$-adaptive and $\rho$-oblivious adversaries are equivalent with no dependence on the domain size. In Appendix A, we give an easy reduction from the standard notions of subtractive noise (without the $\varnothing$ element added to the domain) to the adversaries defined by $\rho_{\mathrm{sub}}$. This reduction, combined with Theorem 3, shows that the standard oblivious and adaptive subtractive adversaries are equivalent.
Additive contamination (Huber’s model [Hub64]): In Huber’s original model [Hub64], rather than directly receive i.i.d. samples from the target distribution $\mathcal{D}$, the algorithm receives i.i.d. samples from $\mathcal{D}^{\prime}$, the mixture distribution
$\mathcal{D}^{\prime}\coloneqq(1-\eta)\mathcal{D}+\eta\mathcal{E}$
where the adversary chooses the outlier distribution $\mathcal{E}$. In the adaptive variant of this model, first a clean sample of $\lceil(1-\eta)n\rceil$ points are drawn i.i.d. from $\mathcal{D}$. Then, the adversary may add $\lfloor\eta n\rfloor$ points arbitrarily. These $n$ points are then randomly permuted so that the algorithm cannot trivially identify which points were added.
Similarly to subtractive contamination, we will use the augmented domain $X^{\prime}\coloneqq X\cup\{\varnothing\}$. For additive noise, we use the cost function
$\rho_{\mathrm{add}}(x,y)\coloneqq\begin{cases}0&\text{if }x=y\\ \frac{1}{\eta}&\text{if }x\neq y\text{ and }x=\varnothing\\ \infty&\text{otherwise}.\end{cases}$
In Appendix A, we give an easy reduction showing how Theorem 3 gives the equivalence between Huber’s contamination model and its adaptive variant. Note that this particular equivalence was already proven by [BLMT22], but for completeness, we show their result can be recovered using our framework.
3.1 Partially-adaptive adversaries
As alluded to in Remark 1, some adversaries lie between the fully oblivious and fully adaptive adversaries. Our results show that such intermediate adversaries are equivalent to their fully oblivious and fully adaptive counterparts. The strategy for proving this equivalency is by showing the intermediate adversary is at least as strong as the oblivious adversary, and that it is no stronger than the adaptive adversary. Since Theorem 2 implies the adaptive adversary is no more powerful than the oblivious adversary, we can conclude that all three adversaries are equivalent. We formalize this approach for the two adversaries described here in Appendix B, showing both are equivalent to additive contamination.
Malicious noise: This model was first defined by [Val85]. In it, The $n$ samples are generated sequentially. For each point, independently with probability $1-\eta$, that point is sampled from $\mathcal{D}$. Otherwise, the adversary chooses an arbitrary corrupted point with full knowledge of previous points generated but no knowledge of future points. Intuitively, this adversary is partially adaptive because when the adversary chooses how to corrupt a point, it has partial knowledge of the sample corresponding to the points generated previously.
The non-independent additive adversary: This model was recently studied in [CHL+23]. In it, the adversary generates $\lfloor\eta n\rfloor$ arbitrary points. The sample is then formed by combining $\lceil(1-\eta)n\rceil$ points drawn i.i.d. from $\mathcal{D}$ with the adversary’s chosen points. Intuitively, this adversary is partially adaptive because the $\eta n$ points it generates need not be i.i.d. from some distribution as they would for fully oblivious adversaries, but the adversary still does not know the sample when choosing corruptions as in an fully adaptive adversary.
4 Technical overview
To prove Theorem 3, we begin with the observation that we can combine the algorithm $A$ and test $T$ into a single function $f\coloneqq T\circ A$. Therefore, it suffices to prove the following.
Theorem 6 (Theorem 3 restated).
For any $n,d\in\mathds{N}$ where $d\geq 2$, domain $X$, and $\varepsilon>0$, let $m=O\left(\frac{n^{3}(\ln d)^{2}}{\varepsilon^{4}}\right)$. Then, for any $f:X^{n}\to\{0,1\}$, cost function $\rho$ with degree $d$, and distribution $\mathcal{D}$ supported on $X$,
$\left|\mathrm{Oblivious\text{-}Max}_{\rho}(f,\mathcal{D})-\mathrm{Adaptive\text{-}Max}_{\rho}(f\circ\Phi_{m\to n},\mathcal{D})\right|\leq\varepsilon.$ (1)
Equation 1 can be written as two separate inequalities,
$\mathrm{Oblivious\text{-}Max}_{\rho}(f,\mathcal{D})\leq\mathrm{Adaptive\text{-}Max}_{\rho}(f\circ\Phi_{m\to n},\mathcal{D})+\varepsilon$ (The easy direction)
$\mathrm{Adaptive\text{-}Max}_{\rho}(f\circ\Phi_{m\to n},\mathcal{D})\leq\mathrm{Oblivious\text{-}Max}_{\rho}(f,\mathcal{D})+\varepsilon$ (The hard direction.)
For the easy direction, Theorem 6 states that for every distinguisher $f$ and every oblivious adversary, there is a corresponding adaptive adversary that appears nearly identical to this distinguisher $f$. We prove a stronger version in which the quantifiers are switched — one adaptive adversary can fool all $f$.
Proposition 4.1 (The adaptive adversary can simulate the oblivious adversary).
For any $m\geq 25n^{2}/\varepsilon^{2}$, distribution $\mathcal{D}$, cost function $\rho$, and oblivious corruption $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$, there is a corresponding adaptive adversary (i.e. a function Corrupt mapping samples $S$ to their corruption $S^{\prime}\in\mathcal{C}_{\rho}(S)$) satisfying, for all $f:X^{n}\to[0,1]$,
$\mathop{{\mathds{E}}\/}_{\bm{S}^{\prime}\sim(\mathcal{D}^{\prime})^{n}}[f(\bm{S}^{\prime})]\leq\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{m}}[f\circ\Phi_{{m}\to n}(\textsc{Corrupt}(\bm{S}))]+\varepsilon.$
[DKK+19, ZJS19] gave similar simulation for specific adversary models (corresponding to specific choices of the cost function), and the proof of Proposition 4.1 is a mostly straightforward generalization of these approaches. We discuss the novel ingredients needed for Proposition 4.1 in Section 8, but will focus the remainder of this technical overview on the more interesting direction: Proving the adaptive adversary is no stronger than the oblivious adversary.
4.1 Deterministic and randomized simulation arguments
In light of Proposition 4.1, it is natural to wonder if a similar “simulation-based” approach can be used to prove the adaptive adversary is no stronger than the oblivious adversary. To formalize this class of approaches, fix a base distribution $\mathcal{D}$, cost function $\rho$, and parameters $m\geq n$. Then define,
  1. 1.
    The set of input distributions over $n$ points the oblivious adversary can create,
    $\mathscr{D}_{\mathrm{oblivious}}^{n,\rho,\mathcal{D}}\coloneqq\{(\mathcal{D}^{\prime})^{n}\mid\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})\}.$
  2. 2.
    The set of input distributions over $n$ points the adaptive adversary can create when the subsampling filter $\Phi_{{m}\to n}$ is used:
[…]
    where $\bm{S}\sim\mathcal{D}^{m}$ is a clean sample and Corrupt is a legal corruption strategy if $\textsc{Corrupt}(S)\in\mathcal{C}_{\rho}(S)$ for all $S\in X^{m}$.
With this statement, the simulation argument in Proposition 4.1 can be succinctly written as follows: For sufficiently large $m$ and any $\mathcal{D}_{\mathrm{oblivious}}\in\mathscr{D}_{\mathrm{oblivious}}^{n,\rho,\mathcal{D}}$, there is some $\mathcal{D}_{\mathrm{adaptive}}\in\mathscr{D}_{\mathrm{adaptive}}^{m,n,\rho,\mathcal{D}}$ for which the total variation distance of $\mathcal{D}_{\mathrm{oblivious}}$ and $\mathcal{D}_{\mathrm{adaptive}}$ is small.
Question 7 (Can the oblivious adversary simulate the adaptive adversary?).
For any $n,\mathcal{D},\rho$ sufficiently large $m$, and any $\mathcal{D}_{\mathrm{adaptive}}\in\mathscr{D}_{\mathrm{adaptive}}^{m,n,\rho,\mathcal{D}}$, is there some $\mathcal{D}_{\mathrm{oblivious}}\in\mathscr{D}_{\mathrm{oblivious}}^{n,\rho,\mathcal{D}}$ for which the total variation distance of $\mathcal{D}_{\mathrm{oblivious}}$ and $\mathcal{D}_{\mathrm{adaptive}}$ is small?
It turns out that there is a simple counterexample to 7. Consider the adaptive and oblivious adversaries described in Example 1 with $\eta=1/2$:
  1. 1.
    For any $S\in X^{m}$, the adaptive adversary can change an arbitrary $m/2$ points within $S$.
  2. 2.
    For any distribution $\mathcal{D}$, the oblivious adversary can choose any $\mathcal{D}^{\prime}$ with a total variation distance of at most $1/2$ from $\mathcal{D}$.
Furthermore, consider perhaps the simplest possible base distribution, $\mathcal{D}=\mathrm{Unif}(\{0,1\})$ and any $m\geq n\coloneqq 2$
After receiving $\bm{S}\sim\mathcal{D}^{m}$, the adaptive adversary can choose a corruption so that $\bm{S}^{\prime}$ either contains only zeros or only ones, with both cases equally probable. They achieve this by flipping all the 00s or all the $1$s in $\bm{S}$, whichever is less frequent (breaking ties uniformly). The result of this approach is that there is some $\mathcal{D}_{\mathrm{adaptive}}\in\mathscr{D}_{\mathrm{adaptive}}$ for which
$\mathcal{D}_{\mathrm{adaptive}}=\mathrm{Unif}([0,0],[1,1]).$
The above distribution is far from any product distribution and, as a result, far from any possible $\mathcal{D}_{\mathrm{oblivious}}$. Therefore, we need a more powerful approach.
Randomized simulations. One key observation about the above counterexample is that, while $\mathcal{D}_{\mathrm{adaptive}}$ is far from any single distribution in $\mathscr{D}_{\mathrm{oblivious}}$, it is close to a mixture of distributions contained within $\mathscr{D}_{\mathrm{oblivious}}$ because the distribution that always outputs $[0,0]$ and the distribution that always outputs $[1,1]$ are both in $\mathscr{D}_{\mathrm{oblivious}}$. This motivates the use of randomized simulations. [^2] We only use one direction of the below equivalence, that randomized simulations imply indistinguishability. We include a proof of the converse as it may be of independent interest. On a historical note, the authors’ original proof was both more complicated and quantitatively weaker than what is presented here, and this converse motivated the search for a simpler argument.
Lemma 4.2 (Randomized simulations are equivalent to indistinguishability).
For any $\mathscr{D}_{1}$ and $\mathscr{D}_{2}$ each a collection of distributions over a domain $X$ and $\varepsilon\geq 0$, if $\mathscr{D}_{2}$ and $X$ are finite, the following are equivalent:
  1. 1.
    There is a randomized simulation of $\mathscr{D}_{1}$ by $\mathscr{D}_{2}$: For any $\mathcal{D}_{1}\in\mathscr{D}_{1}$, there is a distribution $\mathcal{D}_{2}$ formed by taking a mixture of distributions from $\mathscr{D}_{2}$ for which,
    $d_{\mathrm{TV}}(\mathcal{D}_{1},\mathcal{D}_{2})\leq\varepsilon.$
  2. 2.
    $\mathscr{D}_{1}$ is $\varepsilon$-indistinguishable from $\mathscr{D}_{2}$: For any $\mathcal{D}_{1}\in\mathscr{D}_{1}$ and $f:X\to[0,1]$, there some some $\mathcal{D}_{2}\in\mathscr{D}_{2}$ for which
    $\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{1}}[f(\bm{x})]\leq\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{2}}[f(\bm{x})]+\varepsilon.$
Furthermore, randomized simulations imply indistinguishability even if $X$ and/or $\mathscr{D}_{2}$ are not finite.
By Lemma 4.2, it suffices to show that any adaptive adversary can be simulated by a randomized oblivious adversary – i.e. one that uses internal randomness to decide which $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$ to choose as its corrupting distribution. This is the approach we take.
4.2 Structure of the simulation argument
Our goal is to simulate an arbitrary adaptive adversary by a randomized oblivious adversary. Given the description of any adaptive adversary (i.e. a function Corrupt mapping any $S\in X^{m}$ to $S^{\prime}\in\mathcal{C}_{\rho}(S)$), we wish to describe a corresponding randomized oblivious adversary (i.e. a random distribution $\bm{\mathcal{D}}^{\prime}$ where $\bm{\mathcal{D}}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$ with probability $1$). Consider two extreme attempts at designing this simulation.
  1. 1.
    The oblivious adversary draws $\bm{S}\sim\mathcal{D}^{m}$, simulates the adaptive corruption $\bm{S}^{\prime}=\textsc{Corrupt}(\bm{S})$, and tries setting $\bm{\mathcal{D}}^{\prime}\coloneqq\mathrm{Unif}(\bm{S}^{\prime})$. The upside of this approach is that its failry easy to show the distribution of $(\bm{\mathcal{D}}^{\prime})^{n}$ is $(\varepsilon=\frac{n^{2}}{m})$-close, in total variation, to $\mathcal{D}_{\mathrm{adaptive}}$. The downside is that these choices of $\bm{\mathcal{D}}^{\prime}$ are typically not legal corruptions for the oblivious adversary, as $\mathrm{Unif}(\bm{S}^{\prime})$ is not guaranteed to be contained in $\mathcal{C}_{\rho}(\mathcal{D})$.
  2. 2.
    The oblivious adversary chooses a single $\mathcal{D}^{\prime}$ which is the average of $\mathrm{Unif}(\textsc{Corrupt}(\bm{S}))$ over $\bm{S}\sim\mathcal{D}^{m}$. It turns out, this $\mathcal{D}^{\prime}$ is guaranteed to be a legal oblivious corruption (see Lemma 4.3 below). Unfortunately, this is a deterministic adversary and so the counterexample to 7 shows that the resulting $\mathcal{D}_{\mathrm{oblivious}}$ will be far from $\mathcal{D}_{\mathrm{adaptive}}$.
Our approach: Our simulation argument falls between these two extremes. At a high level, we group the adaptively corrupted samples $\bm{S}^{\prime}$ into a moderate number of groups that make “similar” corruptions. The oblivious adversary then
  1. 1.
    Draws a random group.
  2. 2.
    Defines $\bm{\mathcal{D}}_{\mathrm{goal}}$ to the average of the $\mathrm{Unif}(\bm{S}^{\prime})$ over all $\bm{S}^{\prime}$ that fall into the chosen group
  3. 3.
    Rounds $\bm{\mathcal{D}}_{\mathrm{goal}}$ to the nearest legal corruption $\bm{\mathcal{D}}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$.
To analyze this approach, we bound two sources of error.
  1. 1.
    Error due to grouping: This is the TV distance between $\mathcal{D}_{\mathrm{adaptive}}$ and the mixture distribution $(\bm{\mathcal{D}}_{\mathrm{goal}})^{n}$.
  2. 2.
    Error due to rounding. This is the average (over the random group) of the TV distance between $(\bm{\mathcal{D}}_{\mathrm{goal}})^{n}$ and $(\bm{\mathcal{D}}^{\prime})^{n}$.
We defer discussion on how to define the groups and bound the error due to grouping to Section 4.3 as our main lemma there is a bit delicate to state. For the error due to rounding, our lemma is simpler. We show that any large group incurs little error due to rounding, implying that as long as there aren’t too many groups, the total error due to rounding is small.
In the below lemma, the event $\bm{E}$ indicates whether $\bm{S}^{\prime}$ falls into a given group.
Lemma 4.3 (Large groups can be rounded to legal corruptions).
For any cost function $\rho$, base distribution $\mathcal{D}$, and $n\leq m$, draw $\bm{S}\sim\mathcal{D}^{m}$, let $\bm{S}^{\prime}\in\mathcal{C}_{\rho}(\bm{S})$ be any legal adaptive corruption and $\bm{E}$ be any event on the same probability space as $\bm{S}^{\prime}$. Let $\mathcal{D}_{\mathrm{goal}}$ be the distribution of $\bm{x}^{\prime}\sim\mathrm{Unif}(\bm{S}^{\prime})$ conditioned on the event $\bm{E}$ occurring. There is legal oblivious corruption $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$ with total variation distance at most
$d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}^{n},(\mathcal{D}^{\prime})^{n})\leq\sqrt{\frac{n\cdot\ln(1/\operatorname{{Pr}}[\bm{E}])}{2m}}.$
Lemma 4.3 generalizes the fact that if the average of $\mathrm{Unif}(\bm{S}^{\prime})$ for all $\bm{S}^{\prime}$ is a valid oblivious corruption (that fact corresponds to $\operatorname{{Pr}}[\bm{E}]=1$). It remains useful as long as the number of groups is at most $2^{o(m)}$, which will be true of the grouping strategy we describe in Section 4.3.
4.3 Grouping with sunflowers
Intuitively, we wish to group $\bm{S}^{\prime}$ that make similar corruptions together. To do so, we specify each group by a “core” $c\in X^{k}$ which groups together all $\bm{S}^{\prime}$ containing this core $c$. Note that this is a “soft” grouping in the sense that a single $\bm{S}^{\prime}\in X^{m}$ can be a member of multiple groups—one for each size-$k$ subset of $\bm{S}^{\prime}$. Our main technical lemma shows that as long as $k$ is large enough, the error due to grouping is small.
To formally state this lemma, we borrow terminology from the study of sunflowers in combinatorics (see e.g. [ER60, ALWZ20]). A sunflower, $S_{1},\ldots,S_{\ell}$, is a collection of sets in which all distinct pairs of sets share the same intersection. This intersection, $c$, is commonly referred to as the “core” or “kernel,” with $P_{1},\ldots,P_{m},P_{i}\coloneqq S_{i}\setminus c$ referred to as the petals. Note that we slightly generalize this terminology by not requiring the petals to be disjoint.
In this language, the key quantity of interest to our application is the distribution of petals that share a given core.
Definition 8 (Petal distribution).
Let $\tilde{\mathcal{D}}$ be any distribution over $X^{m}$. For any $k\leq m$ and core $c\in X^{k}$ we define the petal distribution for core $c$, $\tilde{\mathcal{P}}(c)$, to be the conditional distribution of $(\bm{S}\setminus c)$ conditioned on $\Phi_{m\to k}(\bm{S})=c$ where $\bm{S}\sim\tilde{\mathcal{D}}$.
Using these petal distributions, we summarize the entire simulation in Figure 1. Our result bounding the error due to grouping is given in Lemma 4.4. Given how intensely sunflowers are studied in the combinatorics literature, it may be of independent interest.
Input: A base distribution $\mathcal{D}$ over $X$, parameters $k\leq m\in\mathds{N}$, and a description of the adaptive adversary, encoded as a function $\textsc{Corrupt}:X^{m}\to X^{m}$. Output: A randomized oblivious adversary simulating the given adaptive adversary. Simulation: The oblivious adversary outputs a random distribution $\bm{\mathcal{D}}^{\prime}$ by, 1. Draw a core $\bm{c}\sim\Phi_{m\to k}(\textsc{Corrupt}(\bm{S}))$ where $\bm{S}\sim\mathcal{D}^{m}$. 2. For any $c\in X^{k}$, define the goal distribution $\bm{\mathcal{D}}_{\mathrm{goal}}(\bm{c})\coloneqq\mathop{{\mathds{E}}\/}_{\bm{P}\sim\tilde{\mathcal{P}}(\bm{c})}[\mathrm{Unif}(\bm{P})],$ where the petal distribution $\tilde{\mathcal{P}}(c)$ is defined in Definition 8 with $\tilde{\mathcal{D}}$ the distribution of $\textsc{Corrupt}(\bm{S})$. 3. Round the goal distribution to $\bm{\mathcal{D}}^{\prime}\coloneqq\operatorname*{\arg\!\min}_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left(d_{\mathrm{TV}}((\bm{\mathcal{D}}_{\mathrm{goal}}(\bm{c}))^{n},(\mathcal{D}^{\prime})^{n})\right).$
Figure 1: A randomized oblivious simulation of any adaptive adversary.
Lemma 4.4 (Grouping using cores has low error).
For any integers $m\geq k_{\max}+n$ and distribution $\tilde{\mathcal{D}}$ over $X^{m}$, let $\mathrm{info}_{r}(\tilde{\mathcal{D}})$ be the mutual information between one coordinate of $\bm{S}\sim\tilde{\mathcal{D}}$ and $r$ other coordinates. There exists some $k\leq k_{\max}$ for which
$d_{\mathrm{TV}}\left(\mathop{{\mathds{E}}\/}_{\bm{S}\sim\tilde{\mathcal{D}}}\left[\Phi_{m\to n}(\bm{S})\right],\mathop{{\mathds{E}}\/}_{\bm{S},\bm{c}\sim\Phi_{m\to k}(\bm{S})}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})\right]\right)\leq\sqrt{\frac{n^{2}\cdot\mathrm{info}_{n+k-1}(\tilde{\mathcal{D}})}{4k_{\max}}}$
where $\mathcal{D}_{\mathrm{goal}}(c)$ is the distribution of $\bm{x}$ created by first drawing $\bm{P}\sim\tilde{\mathcal{P}}(c)$ (for $\tilde{\mathcal{P}}$ defined in Definition 8) and then taking $\bm{x}\sim\mathrm{Unif}(\bm{P})$.
Lemma 4.4 shows that, for large enough $k$, the distance between the target distribution $\mathcal{D}_{\mathrm{adaptive}}$ and the goal distribution $\mathcal{D}_{\mathrm{goal}}(\bm{c})$ is small. Note that the larger $k$ is, the smaller the groups are, and hence the error due to rounding in Lemma 4.3 will be larger. We choose $k$ to balance these two sources of error.
The quantity $\mathrm{info}_{r}(\tilde{\mathcal{D}})$ captures a notion of how far from a product distribution $\bm{S}\sim\tilde{\mathcal{D}}$ is. It is always bounded by $\ln|X|$, which is enough to recover Theorem 2. To remove the dependence on the size of the domain, we show this quantity is bounded by the degree of the cost function.
Lemma 4.5 (Bounding mutual information for low-degree corruptions).
Consider any cost function $\rho$ with degree $d$ and function $\textsc{Corrupt}:X^{m}\to X^{m}$ mapping each $S\in X^{m}$ to some $S^{\prime}\in\mathcal{C}_{\rho}(S)$. Then, for any random variable $\bm{S}$ on $X^{m}$ where $\bm{S}_{1},\ldots,\bm{S}_{m}$ are independent, $\tilde{\mathcal{D}}$ the distribution of $\textsc{Corrupt}(S)$, and $r<m$,
$\mathrm{info}_{r}(\tilde{\mathcal{D}})\leq\frac{m}{m-r}\cdot\ln d.$
We give the proofs of Lemmas 4.4 and 4.5 in Section 7. For now, we’ll just briefly mention that the proof of Lemma 4.4 upper bounds the desired TV distance by an expression involving various quantities measuring the mutual information between different coordinates of $\bm{S}$, and then upper bounds these quantities in terms of $\mathrm{info}_{r}(\tilde{\mathcal{D}})$.
5 Preliminaries
Indexing. For any $n\in\mathds{N}$, we use $[n]$ as shorthand for $\{1,2,\ldots,n\}$. Similarly, for $n\leq m\in\mathds{N}$, we use $[n,m]$ as shorthand for $\{n,n+1,\ldots,m\}$. For any multiset $S\in X^{m}$, we use $S_{i}$ to denote the $i^{\text{th}}$ element of $S$. For any $I\subseteq[m]$, we use $S_{I}$ to denote the multiset containing $(S_{I_{1}},S_{I_{2}},\ldots)$. We’ll also use $S_{\leq j}$ and $S_{<j}$ as shorthand for $S_{[j]}$ and $S_{[j-1]}$ respectively. For any permutation $\sigma:[m]\to[m]$ and $S\in X^{m}$, we’ll use $\sigma(S)$ as shorthand for the multiset in $X^{m}$ satisfying $\sigma(S)_{i}=S_{\sigma(i)}$.
Random variables and distributions. We use boldfont to denote random variables and calligraphic font to denote distributions (e.g. $\bm{x}\sim\mathcal{D}$). For a distribution $\mathcal{D}$, we will use $\bm{x}_{1},\ldots,\bm{x}_{n}\overset{\mathrm{iid}}{\sim}\mathcal{D}$ and $\bm{x}\sim\mathcal{D}^{n}$ interchangeably to denote that $\bm{x}_{1},\ldots,\bm{x}_{n}$ are independent and identically distributed according to $\mathcal{D}$. For any distributions $\mathcal{D}_{1},\mathcal{D}_{2}$, we use $\mathcal{D}_{1}\times\mathcal{D}_{2}$ to denote the product distribution of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. We denote mixture distributions as convex combinations (e.g. $\mathcal{D}_{\mathrm{mix}}=1/3\cdot\mathcal{D}_{1}+2/3\cdot\mathcal{D}_{2}$).
We use the following standard concentration inequality.
Fact 5.1 (Chernoff bound).
Let $\bm{x}_{1},\ldots,\bm{x}_{n}$ be independent random variables on $\{0,1\}$, and $\bm{X}$ their sum. For $\mu\coloneqq\mathop{{\mathds{E}}\/}[\bm{X}]$,
$\operatorname{{Pr}}[\bm{X}\geq 2\mu]\leq e^{-\mu/3}\quad\quad\text{and}\quad\quad\operatorname{{Pr}}[\bm{X}\leq\mu/2]\leq e^{-\mu/8}.$
We will also use two commonly studied families of random variables. For any $p\in[0,1]$, we use $\mathrm{Ber}(p)$ to denote the distribution that takes on value $1$ with probability $p$ and takes on 00 otherwise. Furthermore, for any $n\in\mathds{N}$, we use $\mathrm{Bin}(n,p)$ to denote the sum of $n$ independent random variables each distributed according to $\mathrm{Ber}(p)$.
TV distance and KL divergence. We use two measures of statistical distance/divergence.
Definition 9 (Total variation distance).
Let $\mathcal{D}$ and $\mathcal{D}^{\prime}$ be any two distributions over the same domain, $X$. The total variation distance between $\mathcal{D}$ and $\mathcal{D}^{\prime}$, is defined as
$d_{\mathrm{TV}}(\mathcal{D},\mathcal{D}^{\prime})\coloneqq\sup_{T:X\to[0,1]}\left\{\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}[T(\bm{x})]-\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}^{\prime}}[T(\bm{x}^{\prime})]\right\}.$
This quantity can be equivalently defined as the infimum over all couplings of $\bm{x}\sim\mathcal{D}$ and $\bm{x}^{\prime}\sim\mathcal{D}^{\prime}$ of $\operatorname{{Pr}}[\bm{x}\neq\bm{x}^{\prime}]$.
TV distance is a true distance in the sense that it satisfies the triangle inequality.
Fact 5.2 (The triangle inequality for TV distance).
For any distributions $\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{D}_{3}$,
$d_{\mathrm{TV}}(\mathcal{D}_{1},\mathcal{D}_{3})\leq d_{\mathrm{TV}}(\mathcal{D}_{1},\mathcal{D}_{2})+d_{\mathrm{TV}}(\mathcal{D}_{2},\mathcal{D}_{3}).$
We can also give a (sometimes coarse) upper bound on the TV distance of product distribution.
Fact 5.3 (Total variation distance of a product).
For any distributions $\mathcal{D}_{1},\mathcal{D}_{2}$ and $n\in\mathds{N}$,
$d_{\mathrm{TV}}(\mathcal{D}_{1}^{n},\mathcal{D}_{2}^{n})\leq n\cdot d_{\mathrm{TV}}(\mathcal{D}_{1},\mathcal{D}_{2}).$
Straight from the definition, we see that TV distance is convex in one argument.
Fact 5.4 (Convexity of TV distance).
For any distributions $\mathcal{D}_{1},\mathcal{D}_{2},$ and $\mathcal{E}$, and mixture weight $\lambda\in[0,1]$,
$d_{\mathrm{TV}}(\mathcal{D}_{\lambda},\mathcal{E})\leq\lambda\cdot d_{\mathrm{TV}}(\mathcal{D}_{1},\mathcal{E})+(1-\lambda)\cdot d_{\mathrm{TV}}(\mathcal{D}_{2},\mathcal{E}),$
where $\mathcal{D}_{\lambda}$ is the mixture $\lambda\mathcal{D}_{1}+(1-\lambda)\mathcal{D}_{2}$.
The other measure of statistical distance/divergence that plays a key role in our results in KL divergence.
Definition 10 (Kullback-Leibler (KL) Divergence).
For distributions $\mathcal{D},\mathcal{E}$ supported on the same domain $X$, the KL divergence between $\mathcal{D}$ and $\mathcal{E}$ as defined as,
$d_{\mathrm{KL}}\left({\mathcal{D}}\,\middle\|\,{\mathcal{E}}\right)\coloneqq\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}\left[\ln\left(\frac{\mathcal{D}(\bm{x})}{\mathcal{E}(\bm{x})}\right)\right],$
where $\mathcal{D}(x)$ and $\mathcal{E}(x)$ denote the probability mass or density functions of $\mathcal{D}$ and $\mathcal{E}$ respectively at the point $x$ (or more generally, $\mathcal{D}(x)/\mathcal{E}(x)$ is the Radon-Nikodym derivative of $\mathcal{D}$ with respect to $\mathcal{E}$).
It will sometimes be more convenient to use the variational (also known as “dual”) form of KL divergence.
Fact 5.5 (Variational form of KL divergence, [DV83, Gra11]).
For distributions $\mathcal{D}$, $\mathcal{E}$ supported on the same domain $X$, the KL divergence between $\mathcal{D}$ and $\mathcal{E}$ is defined as,
$d_{\mathrm{KL}}\left({\mathcal{D}}\,\middle\|\,{\mathcal{E}}\right)=\sup_{f:X\to\mathds{R}}\left\{\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}[f(\bm{x})]-\ln\left(\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{E}}\left[e^{f(\bm{x})}\right]\right)\right\}.$
A standard property of KL divergence is that it satisfies additivity when both arguments are product distributions.
Fact 5.6 (Additivity of KL divergence).
For any distributions $\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{E}_{1},\mathcal{E}_{2}$
$d_{\mathrm{KL}}\left({\mathcal{D}_{1}\times\mathcal{D}_{2}}\,\middle\|\,{\mathcal{E}_{1}\times\mathcal{E}_{2}}\right)=d_{\mathrm{KL}}\left({\mathcal{D}_{1}}\,\middle\|\,{\mathcal{E}_{1}}\right)+d_{\mathrm{KL}}\left({\mathcal{D}_{2}}\,\middle\|\,{\mathcal{E}_{2}}\right).$
A second standard property is that KL divergence is convex.
Fact 5.7 (Convexity of KL Divergence).
For any distributions $\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{E}_{1},\mathcal{E}_{2}$ and mixture weight $\lambda\in[0,1]$,
$d_{\mathrm{KL}}\left({\mathcal{D}_{\lambda}}\,\middle\|\,{\mathcal{E}_{\lambda}}\right)\leq\lambda\cdot d_{\mathrm{KL}}\left({\mathcal{D}_{1}}\,\middle\|\,{\mathcal{E}_{1}}\right)+(1-\lambda)\cdot d_{\mathrm{KL}}\left({\mathcal{D}_{2}}\,\middle\|\,{\mathcal{E}_{2}}\right)$
where $\mathcal{D}_{\lambda}$ and $\mathcal{E}_{\lambda}$ are the mixtures $\lambda\mathcal{D}_{1}+(1-\lambda)\mathcal{D}_{2}$ and $\lambda\mathcal{E}_{1}+(1-\lambda)\mathcal{E}_{2}$ respectively.
Unlike TV distance, KL divergence is not a true distance in the sense that it does not satisfy triangle inequality. For us, it will suffices that it upper bounds TV distance via Pinkser’s inequality [Pin64]. [Can22] has a nice summary of different forms and proofs of the below inequality and appropriate references for each.
Fact 5.8 (Pinkser’s inequality [Pin64, Can22]).
For any distributions $\mathcal{D},\mathcal{E}$,
$d_{\mathrm{TV}}(\mathcal{D},\mathcal{E})\leq\sqrt{\frac{d_{\mathrm{KL}}\left({\mathcal{D}}\,\middle\|\,{\mathcal{E}}\right)}{2}}.$
Mutual information.
Definition 11 (Mutual information).
For random variables $\bm{x},\bm{y}$ jointly distribution according to a distribution $\mathcal{D}$, let $\mathcal{D}_{x}$ and $\mathcal{D}_{y}$ be the marginal distributions of $\bm{x}$ and $\bm{y}$ respectively, and $\mathcal{D}_{x\mid y}$ be the marginal distribution of $\bm{x}$ conditioned on $\bm{y}=y$. The mutual information between $\bm{x}$ and $\bm{y}$ is defined as
$I(\bm{x};\bm{y})=d_{\mathrm{KL}}\left({\mathcal{D}}\,\middle\|\,{\mathcal{D}_{x}\times\mathcal{D}_{y}}\right)=\mathop{{\mathds{E}}\/}_{\bm{y}}\left[d_{\mathrm{KL}}\left({\mathcal{D}_{x\mid\bm{y}}}\,\middle\|\,{\mathcal{D}_{x}}\right)\right]$
Definition 12 (Conditional mutual information).
For random variables $\bm{x},\bm{y},\bm{z}$ jointly distributed, the mutual information of $\bm{x}$ and $\bm{y}$ conditioned on $\bm{z}$ is
$I(\bm{x};\bm{y}\mid\bm{z})\coloneqq\mathop{{\mathds{E}}\/}_{\bm{z}^{\prime}\sim\mathcal{D}_{z}}\left[I((\bm{x}\mid\bm{z}=\bm{z}^{\prime});(\bm{y}\mid\bm{z}=\bm{z}^{\prime})\right]$
where $\mathcal{D}_{z}$ is the marginal distribution of $\bm{z}$.
The chain rule connects mutual information and conditional mutual information.
Fact 5.9 (Chain rule for mutual information).
For any $\bm{x},\bm{y},\bm{z}$,
$I(\bm{x};(\bm{y},\bm{z}))=I(\bm{x};\bm{z})+I(\bm{x};\bm{y}\mid\bm{z}).$
This is sometimes rewritten as
$I(\bm{x};\bm{y}\mid\bm{z})=I(\bm{x};(\bm{y},\bm{z}))-I(\bm{x};\bm{z}).$
Mutual information is always nonnegative
Fact 5.10 (Nonnegativity of mutual information).
For any random variables $\bm{x},\bm{y}$,
$I(\bm{x};\bm{y})\geq 0.$
As an easy consequence of the chain rule and mutual information being nonnegative, we have that mutual information can only increase if we consider more information.
Fact 5.11.
For any random variables $\bm{x},\bm{y},\bm{z}$,
$I(\bm{x};(\bm{y},\bm{z}))\geq I(\bm{x};\bm{y}).$
Mutual information is also symmetric.
Fact 5.12 (Symmetry of mutual information).
For any random variables $\bm{x},\bm{y}$,
$I(\bm{x};\bm{y})=I(\bm{y};\bm{x}).$
Another nice property of mutual information is that it is bounded by the support size of each variable.
Fact 5.13 (Mutual information with a finite support).
For any random variables $\bm{x},\bm{y}$, if one of $\bm{x}$ or $\bm{y}$ has a finite support of size $d$, then,
$I(\bm{x};\bm{y})\leq\ln d.$
6 Rounding to a legal corruption
In this section, we prove Lemma 4.3, restated below for convenience. After proving it, in Section 6.3, will show how to apply Lemma 4.3 to bound the error due to rounding for our specific simulation. See 4.3
This proof is broken into the following two pieces. First, we will show that Claim 6.1 holds in the adversary cannot make corruptions.
Claim 6.1.
For any distribution $\mathcal{D}$, draw $\bm{S}\sim\mathcal{D}^{m}$ and let $\bm{E}$ be any event on the same probability space as $\bm{S}$. For $\mathcal{D}_{\mathrm{goal}}$ the distribution of $\bm{x}\sim\mathrm{Unif}(\bm{S})$ conditioned on the event $\bm{E}$ occurring,
$d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}^{n},\mathcal{D}^{n})\leq\sqrt{\frac{n\cdot\ln(1/\operatorname{{Pr}}[\bm{E}])}{2m}}.$
This is a special case of Lemma 4.3 corresponding to when the cost function is simply
$\rho(x,y)=\begin{cases}0&\text{if }x=y\\ \infty&\text{if }x\neq y.\end{cases}$
Then, we will use the following to show that the special case in Claim 6.1 is sufficient to prove the more general case.
Claim 6.2 (Lipschitzness of corruptions).
For any cost function $\rho$, $n\in\mathds{N}$, distributions $\mathcal{D}$ and $\mathcal{D}_{\mathrm{goal}}$, and any $\mathcal{D}_{\mathrm{goal}}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D}_{\mathrm{goal}})$, there is some $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$ for which
$d_{\mathrm{TV}}((\mathcal{D}_{\mathrm{goal}}^{\prime})^{n},(\mathcal{D}^{\prime})^{n})\leq d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}^{n},\mathcal{D}^{n}).$
Proof.
Let $\bm{S},\bm{S}^{\prime},$ and $\bm{E}$ be the random variables defined in Lemma 4.3. Define $\mathcal{D}_{\mathrm{goal}}$ to be the distribution of a uniform element of $\bm{E}$ conditioned on $\bm{E}$ and $\mathcal{D}_{\mathrm{goal}}^{\prime}$ the distribution of a uniform element of $\bm{S}$ conditioned on $\bm{E}$. Then, Claim 6.1 gives that,
$d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}^{n},\mathcal{D}^{n})\leq\sqrt{\frac{n\cdot\ln(1/\operatorname{{Pr}}[\bm{E}])}{2m}}.$
Our goal is to prove that $d_{\mathrm{TV}}((\mathcal{D}_{\mathrm{goal}}^{\prime})^{n},(\mathcal{D}^{\prime})^{n})$ is bounded by the same quantity. Therefore, by Claim 6.2, it suffices to prove that $\mathcal{D}_{\mathrm{goal}}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D}_{\mathrm{goal}})$. This means showing a coupling of $\bm{x}\sim\mathcal{D}_{\mathrm{goal}}$ and $\bm{x}^{\prime}\sim\mathcal{D}_{\mathrm{goal}}^{\prime}$ for which $\mathop{{\mathds{E}}\/}[\rho(\bm{x},\bm{x}^{\prime})]\leq 1$.
For this coupling, let $\bm{S}$ and $\bm{S}^{\prime}$ be as earlier and $\bm{i}\sim\mathrm{Unif}([m])$ be a uniform index. Set $\bm{x}\coloneqq\bm{S}_{\bm{i}}$ and $\bm{x}^{\prime}\sim\coloneqq\bm{S}_{\bm{i}}^{\prime}$. By definition, $\bm{x}$ and $\bm{x}^{\prime}$ have the correct marginal distributions. Finally,
$\mathop{{\mathds{E}}\/}\left[\rho(\bm{x},\bm{x}^{\prime})\right]=\mathop{{\mathds{E}}\/}\left[\frac{1}{m}\sum_{i\in[m]}\rho(\bm{S}_{i},\bm{S}_{i}^{\prime})\right]\leq 1.\qed$
6.1 The special case when the adversary cannot corrupt: Proof of Claim 6.1
This proof can be summarized with a series of steps, each using a basic property of KL divergence.
  1. 1.
    First, we observe that KL-divergence between the distribution of $\bm{S}$ conditioned on $\bm{E}$ and $\mathcal{D}^{m}$ is at most $\ln(1/\operatorname{{Pr}}[\bm{E}])$. Intuitively, this is because the probability of observing any sample conditioned on $\bm{E}$ can only be a factor of $1/\operatorname{{Pr}}[\bm{E}]$ larger than without conditioning.
  2. 2.
    Then, we use Proposition 6.4 to show the above implies that $d_{\mathrm{KL}}\left({\mathcal{D}_{\mathrm{goal}}}\,\middle\|\,{\mathcal{D}}\right)$ is at most $\frac{\ln(1/\operatorname{{Pr}}[\bm{E}])}{m}$.
  3. 3.
    The desired result then follows from simple applications of the additivity of KL divergence (Fact 5.6) and the relation between KL divergence and TV distance (Pinkser’s inequality, Fact 5.8).
For the first step, we’ll use the following simple property of KL divergence.
Proposition 6.3 (KL divergence of a conditional distribution).
For any random variable $\bm{x}$ with support $X$ and event $\bm{E}$, let $\mathcal{D}_{x}$ be the marginal distribution of $\bm{x}$ and $\mathcal{D}_{x\mid E}$ be the marginal distribution of $\bm{x}$ conditioned on the event $\bm{E}$ occurring. Then
$d_{\mathrm{KL}}\left({\mathcal{D}_{x\mid E}}\,\middle\|\,{\mathcal{D}_{x}}\right)\leq\ln(1/\operatorname{{Pr}}[\bm{E}]).$
Proof.
Using Fact 5.5,
$d_{\mathrm{KL}}\left({\mathcal{D}_{x\mid E}}\,\middle\|\,{\mathcal{D}_{x}}\right)$ $=\sup_{f:X\to\mathds{R}}\left\{\mathop{{\mathds{E}}\/}[f(\bm{x})\mid\bm{E}]-\ln\left(\mathop{{\mathds{E}}\/}\left[e^{f(\bm{x})}\right]\right)\right\}$
$=\sup_{f:X\to\mathds{R}}\left\{\mathop{{\mathds{E}}\/}[f(\bm{x})\mid\bm{E}]-\ln\left(\operatorname{{Pr}}[\bm{E}]\cdot\mathop{{\mathds{E}}\/}\left[e^{f(\bm{x})}\mid\bm{E}\right]+\operatorname{{Pr}}[\overline{\bm{E}}]\cdot\mathop{{\mathds{E}}\/}\left[e^{f(\bm{x})}\mid\overline{\bm{E}}\right]\right)\right\}$
$\leq\sup_{f:X\to\mathds{R}}\left\{\mathop{{\mathds{E}}\/}[f(\bm{x})\mid\bm{E}]-\ln\left(\operatorname{{Pr}}[\bm{E}]\cdot\mathop{{\mathds{E}}\/}\left[e^{f(\bm{x})}\mid\bm{E}\right]\right)\right\}$ (Nonnegativity of $x\mapsto e^{x}$)
$\leq\sup_{f:X\to\mathds{R}}\left\{\mathop{{\mathds{E}}\/}[f(\bm{x})\mid\bm{E}]-\ln\left(\operatorname{{Pr}}[\bm{E}]\cdot e^{\mathop{{\mathds{E}}\/}\left[f(\bm{x})\mid\bm{E}\right]}\right)\right\}$ (Jensen’s inequality)
$\leq-\ln(\operatorname{{Pr}}[\bm{E}])=\ln(1/\operatorname{{Pr}}[\bm{E}]).$
For the second step above, we use that KL divergence is superadditive when the second argument is a product distribution.
Proposition 6.4 (Superadditivity of KL divergence).
For any distribution $\mathcal{D}$ over $X^{m}$, and distribution $\mathcal{E}$ over $X$,
$d_{\mathrm{KL}}\left({\mathcal{D}}\,\middle\|\,{\mathcal{E}^{m}}\right)\geq\sum_{i=1}^{m}d_{\mathrm{KL}}\left({\mathcal{D}_{i}}\,\middle\|\,{\mathcal{E}}\right)$
where $\mathcal{D}_{i}$ denotes the $i^{\text{th}}$ marginal of $i$.
Proof.
By Fact 5.5, our goal is to show that
$\sup_{F:X^{m}\to\mathds{R}}\left(\mathop{{\mathds{E}}\/}_{\bm{X}\sim\mathcal{D}}[F(\bm{X})]-\ln\left(\mathop{{\mathds{E}}\/}_{\bm{X}\sim\mathcal{E}^{m}}\left[e^{F(\bm{X})}\right]\right)\right)\geq\sup_{f_{1},\ldots,f_{m}:X\to\mathds{R}}\left(\sum_{i=1}^{m}\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{i}}[f_{i}(\bm{x})]-\ln\left(\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{E}}\left[e^{f_{i}(\bm{x})}\right]\right)\right).$
It therefore suffices to show that for any choice of $f_{1},\ldots,f_{m}$, there exists a choice of $F$ so that the left-hand side of the above equation is at least as large as the right-hand side. Taking
$F(X)=\sum_{i=1}^{m}f(x_{i})$
we show that both sides are equal:
$\mathop{{\mathds{E}}\/}_{\bm{X}\sim\mathcal{D}}[F(\bm{X})]-\ln\left(\mathop{{\mathds{E}}\/}_{\bm{X}\sim\mathcal{E}^{m}}\left[e^{F(\bm{X})}\right]\right)$ $=\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{i}}\left[\sum_{i=1}^{m}f_{i}(\bm{x})\right]-\ln\left(\mathop{{\mathds{E}}\/}_{\bm{X}\sim\mathcal{E}^{m}}\left[e^{\sum_{i=1}^{m}f_{i}(\bm{X}_{i})}\right]\right)$
$=\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{i}}\left[\sum_{i=1}^{m}f_{i}(\bm{x})\right]-\ln\left(\prod_{i=1}^{m}\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{E}}\left[f_{i}(\bm{x})\right]\right)$
$=\sum_{i=1}^{m}\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{i}}[f_{i}(\bm{x})]-\ln\left(\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{E}}\left[e^{f_{i}(\bm{x})}\right]\right).\qed$
We are now ready the main result of this subsection:
Proof of Claim 6.1.
Let $\bm{\sigma}:[m]\to[m]$ be a uniform permutation that is independent of $(\bm{S},\bm{E})$, and let $\mathcal{S}$ be the distribution of $\bm{\sigma}(\bm{S})$ conditioned on $\bm{E}$ occurring where $\bm{\sigma}(\bm{S})$ is shorthand for sample whose $i^{\text{th}}$ value is $\bm{S}_{\sigma(i)}$. We first claim that
$d_{\mathrm{KL}}\left({\mathcal{S}}\,\middle\|\,{\mathcal{D}^{m}}\right)\leq\ln\left(\frac{1}{\operatorname{{Pr}}[\bm{E}]}\right).$
The above follows from Proposition 6.3 and that the distribution of $\bm{\sigma}(\bm{S})$ is $\mathcal{D}^{m}$ (because $\mathcal{D}^{m}$ is a permutation invariant distribution).
Next, we note that for any $i\in[m]$, the marginal distribution $\mathcal{S}_{i}$ is identical to $\mathcal{D}_{\mathrm{goal}}$ (here we crucially use that $\mathcal{S}$ is the distribution of $\bm{S}\mid\bm{E}$ after a uniform permutation is applied). Therefore, using Proposition 6.4,
$d_{\mathrm{KL}}\left({\mathcal{D}_{\mathrm{goal}}}\,\middle\|\,{\mathcal{D}}\right)=\frac{1}{m}\cdot\sum_{i=1}^{m}d_{\mathrm{KL}}\left({\mathcal{S}_{i}}\,\middle\|\,{\mathcal{D}}\right)\leq\frac{d_{\mathrm{KL}}\left({\mathcal{S}}\,\middle\|\,{\mathcal{D}^{m}}\right)}{m}\leq\frac{\ln(1/\operatorname{{Pr}}[\bm{E}])}{m}.$
Finally, we bound,
$d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}^{n},\mathcal{D}^{n})$ $\leq\sqrt{\frac{d_{\mathrm{KL}}\left({\mathcal{D}_{\mathrm{goal}}^{n}}\,\middle\|\,{\mathcal{D}^{n}}\right)}{2}}$ (Fact 5.8)
$=\sqrt{\frac{n\cdot d_{\mathrm{KL}}\left({\mathcal{D}_{\mathrm{goal}}}\,\middle\|\,{\mathcal{D}}\right)}{2}}$ (Fact 5.6)
$=\sqrt{\frac{n\cdot\ln(1/\operatorname{{Pr}}[\bm{E}])}{2m}}.\qed$
6.2 Generalizing to when the adversary can corrupt: Proof of Claim 6.2
Proof.
By Definition 9, it suffices to show that for any coupling of $(\bm{x}_{1},\ldots,\bm{x}_{n})\sim\mathcal{D}_{\mathrm{goal}}^{n}$ and $(\bm{y}_{1},\ldots,\bm{y}_{n})\sim\mathcal{D}^{n}$, there is a coupling of $(\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime})\sim(\mathcal{D}_{\mathrm{goal}}^{\prime})^{n}$ and $(\bm{y}_{1}^{\prime},\ldots,\bm{y}_{n}^{\prime})\sim(\mathcal{D}^{\prime})^{n}$ for which
$\operatorname{{Pr}}[\bm{x}^{\prime}\neq\bm{y}^{\prime}]\leq\operatorname{{Pr}}[\bm{x}\neq\bm{y}].$
Since $\mathcal{D}_{\mathrm{goal}}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D}_{\mathrm{goal}})$, there is a coupling of $\bm{x}_{1}\sim\mathcal{D}_{\mathrm{goal}}$ and $\bm{x}_{1}^{\prime}\sim\mathcal{D}_{\mathrm{goal}}^{\prime}$ for which $\mathop{{\mathds{E}}\/}[\rho(\bm{x}_{1},\bm{x}_{1}^{\prime})]\leq 1$. By taking $n$ independent copies of this coupling, we can have a coupling of $(\bm{x}_{1},\ldots,\bm{x}_{n})\sim\mathcal{D}_{\mathrm{goal}}^{n}$ and $(\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime})\sim(\mathcal{D}_{\mathrm{goal}}^{\prime})^{n}$ for which $\mathop{{\mathds{E}}\/}[\rho(\bm{x}_{i},\bm{x}_{i}^{\prime})]\leq 1$ for all $i\in[n]$.
Now, we combine the two aforementioned couplings of $(\bm{x},\bm{y})$ and of $(\bm{x},\bm{x}^{\prime})$. Formally, this means drawing $\bm{x}$ from its marginal distribution and then drawing $\bm{y}\mid\bm{x}$ and $\bm{x}^{\prime}\mid\bm{x}$ according to the two aforementioned couplings. We then define, $\bm{y}_{1}^{\prime},\ldots,\bm{y}_{n}^{\prime}$ as, for all $i\in[n]$,
$\bm{y}_{i}^{\prime}=\begin{cases}\bm{x}_{i}^{\prime}&\text{if }\bm{x}_{i}=\bm{y}_{i}\\ \bm{y}_{i}&\text{otherwise,}\end{cases}$ (2)
and let $\mathcal{D}^{\prime}$ be the distribution of $\bm{y}_{\bm{i}}$ for $\bm{i}\sim\mathrm{Unif}([n])$. This proof is completed by the following two claims:
Claim 1, $\operatorname{{Pr}}[\bm{x}^{\prime}\neq\bm{y}^{\prime}]\leq\operatorname{{Pr}}[\bm{x}\neq\bm{y}]$: For this, we simply observe from Equation 2 that if $\bm{x}=\bm{y}$ then $\bm{x}^{\prime}=\bm{y}^{\prime}$. Therefore, $\operatorname{{Pr}}[\bm{x}^{\prime}=\bm{y}^{\prime}]\geq\operatorname{{Pr}}[\bm{x}=\bm{y}]$, and negating this gives the desired result.
Claim 2, $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$: For this, it suffices to show that $\mathop{{\mathds{E}}\/}[\rho(\bm{y}_{\bm{i}},\bm{y}^{\prime}_{\bm{i}})]\leq 1$. We bound,
$\mathop{{\mathds{E}}\/}[\rho(\bm{y}_{\bm{i}},\bm{y}^{\prime}_{\bm{i}})]$ $=\mathop{{\mathds{E}}\/}[\rho(\bm{x}_{\bm{i}},\bm{x}_{\bm{i}}^{\prime})\cdot\mathds{1}[\bm{x}_{\bm{i}}=\bm{y}_{\bm{i}}]]$ (Equation 2 and $\rho(y,y)=0$ for any $y$)
$\leq\mathop{{\mathds{E}}\/}[\rho(\bm{x}_{\bm{i}},\bm{x}_{\bm{i}}^{\prime})]\leq 1,$
where the last inequality uses that $\mathop{{\mathds{E}}\/}[\rho(\bm{x}_{i},\bm{x}_{i}^{\prime})]\leq 1$ for any $i\in[n]$. ∎
6.3 Error due to rounding of our strategy
We prove the following.
Lemma 6.5 (Error due to rounding for our corruption strategy).
For any cost function $\rho$ with degree $d$, base distribution $\mathcal{D}$ and $k,n\leq m$, draw $\bm{S}\sim\mathcal{D}^{m}$ and let $\bm{S}^{\prime}\in\mathcal{C}_{\rho}(\bm{S})$ be any legal adaptive corruption of $\bm{S}$. For any core $c^{\prime}\in X^{k}$, let $\mathcal{D}_{\mathrm{goal}}(c^{\prime})$ be the distribution of a uniform element of $\bm{S}^{\prime}\setminus c^{\prime}$ conditioned on $\Phi_{m\to k}(\bm{S}^{\prime})=c^{\prime}$. Then,
$\mathop{{\mathds{E}}\/}_{\bm{c}^{\prime}\sim\Phi_{m\to k}(\bm{S}^{\prime})}\left[\inf_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}(\bm{c}^{\prime})^{n},(\mathcal{D}^{\prime})^{n})\right\}\right]\leq\sqrt{\frac{nk\ln d}{2(m-k)}}+\frac{nk}{m}.$
Proof.
We begin by introducing some convenient notation: Let $\bm{I}\subseteq[m]$ be a uniform subset of $k$ indices from $m$ and define,
$\bm{c}=\bm{S}_{\bm{I}}$ $\bm{P}=\bm{S}_{[m]\setminus\bm{I}}$
$\bm{c}^{\prime}=\bm{S}^{\prime}_{\bm{I}}$ $\bm{P}^{\prime}=\bm{S}^{\prime}_{[m]\setminus\bm{I}}$
In this notation, the distribution $\mathcal{D}_{\mathrm{goal}}(c^{\prime})$ is the distribution of a uniform element of $\bm{P}^{\prime}$ conditioned on $\bm{c}^{\prime}=c^{\prime}$. It will be helpful for us to define the distribution $\mathcal{D}_{\mathrm{goal}}(c,c^{\prime})$ which is the distribution of a uniform element of $\bm{P}^{\prime}$ conditioned on $\bm{c}=c$ and $\bm{c}^{\prime}=c^{\prime}$. In this notation,
$\mathcal{D}_{\mathrm{goal}}(c^{\prime})=\mathop{{\mathds{E}}\/}_{\bm{c}}[\mathcal{D}_{\mathrm{goal}}(\bm{c},c^{\prime})\mid\bm{c}^{\prime}=c^{\prime}].$ (3)
We say that the pair $(c,c^{\prime})$ are “feasible” if
$\rho(c_{i},c_{i}^{\prime})<\infty\quad\quad\text{for all }i\in[k],$
and define,
$p(c,c^{\prime})=\operatorname{{Pr}}[\bm{c}^{\prime}=c^{\prime}\mid\bm{c}=c].$
In the above notation, $p(c,c^{\prime})$ can only be nonzero if $(c,c^{\prime})$ are feasible. The first step of this proof is to show, for any $c,c^{\prime}$,
$\inf_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}(c,c^{\prime})^{n},(\mathcal{D}^{\prime})^{n})\right\}\leq\sqrt{\frac{n\ln(1/p(c,c^{\prime}))])}{2(m-k)}}+\frac{nk}{m}.$ (4)
The desired result will hold by, roughly speaking, observing that there are at most $d^{k}$ feasible $c^{\prime}$ for any $c$, and replacing $p(c,c^{\prime})$ with $d^{-k}$.
The first key observation in showing Equation 4 holds is that, even if we condition on $\bm{c}=c$ the distribution of $\bm{P}$ is still simply $\mathcal{D}^{m-k}$, this is because we had originally draw $\bm{S}\sim\mathcal{D}^{m}$ and so $\bm{c}$ and $\bm{P}$ are independent.
Second, we analyze the relationship between $\bm{P}$ and $\bm{P}^{\prime}$. Since $\bm{S}^{\prime}\in\mathcal{C}_{\rho}(\bm{S})$, we have that
$\sum_{i\in[m-k]}\rho(\bm{P}_{i},\bm{P}^{\prime}_{i})\leq\sum_{i\in[m]}\rho(\bm{S}_{i},\bm{S}_{i}^{\prime})\leq m.$
Note that this does not necessarily mean that $\bm{P}^{\prime}\in\mathcal{C}_{\rho}(\bm{P})$, as that would require $\sum_{i\in[m-k]}\rho(\bm{P}_{i},\bm{P}^{\prime}_{i})\leq m-k$. From the above, we see that if the adversary made no corruptions to $\bm{c}$, it can make extra corruptions to $\bm{P}$. Instead, we get that $\bm{P}^{\prime}$ is a valid corruption of $\bm{P}$ for a slightly more permissive cost function,
$\bm{P}^{\prime}\in\mathcal{C}_{\bar{\rho}}(\bm{P})\quad\quad\text{for}\quad\quad\bar{\rho}(x,y)=\frac{m-k}{m}\cdot\rho(x,y).$
We are now ready to apply Lemma 4.3. Conditioned on $\bm{c}=c$, we have shown that,
  1. 1.
    $\bm{P}$ is distributed according to $\mathcal{D}^{m-k}$.
  2. 2.
    $\bm{P}^{\prime}\in\mathcal{C}_{\bar{\rho}}(\bm{P})$ with probability $1$.
Then, since $\mathcal{D}_{\mathrm{goal}}(c,c^{\prime})$ is the distribution of $\mathrm{Unif}(\bm{P}^{\prime})$ conditioned on $\bm{c}^{\prime}=c^{\prime}$ and $\bm{c}=c^{\prime}$, Lemma 4.3 gives that there is some $\bar{\mathcal{D}}\in\mathcal{C}_{\bar{\rho}}(\mathcal{D})$ for which,
$d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}(c,c^{\prime})^{n},(\bar{\mathcal{D}})^{n})\leq\sqrt{\frac{n\ln(1/p(c,c^{\prime}))}{2(m-k)}}.$
To recover Equation 4 from the above, we will show that for any $\bar{\mathcal{D}}\in\mathcal{C}_{\bar{\rho}}(\mathcal{D})$, there is some $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$ for which,
$d_{\mathrm{TV}}(\mathcal{D}^{\prime},\bar{\mathcal{D}})\leq\frac{k}{m}.$
The desired bound then follows from Facts 5.3 and 5.2. For the above, since $\bar{\mathcal{D}}\in\mathcal{C}_{\bar{\rho}}(\mathcal{D})$, there is a coupling of $\bm{x}\sim\mathcal{D}$ and $\bar{\bm{x}}\sim\bar{\mathcal{D}}$ for which $\mathop{{\mathds{E}}\/}[\bar{\rho}(\bm{x},\bar{\bm{x}})]\leq 1$. Let $\mathcal{D}^{\prime}$ be the distribution of
$\bm{x}^{\prime}=\begin{cases}\bar{\bm{x}}&\text{with probability }\frac{m-k}{m}\\ \bm{x}&\text{with probability }\frac{k}{m}.\end{cases}$
Then, $d_{\mathrm{TV}}(\mathcal{D}^{\prime},\bar{\mathcal{D}})\leq\operatorname{{Pr}}[\bm{x}^{\prime}\neq\bar{\bm{x}}]\leq\frac{k}{m}$, and $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$ because
$\mathop{{\mathds{E}}\/}[\rho(\bm{x},\bm{x}^{\prime})]=\frac{m-k}{m}\mathop{{\mathds{E}}\/}[\rho(\bm{x},\bar{\bm{x}})]=\frac{m-k}{m}\cdot\frac{m}{m-k}\mathop{{\mathds{E}}\/}[\bar{\rho}(\bm{x},\bar{\bm{x}})]\leq 1.$
This completes the proof of Equation 4. Next, we use this to understand the distance between $\mathcal{D}_{\mathrm{goal}}(c^{\prime})$ and the nearest legal corruption. First, for any $c^{\prime}$, by the convexity of TV distance (Fact 5.4) and Equation 3,
$\inf_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}(c^{\prime})^{n},(\mathcal{D}^{\prime})^{n})\right\}\leq\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\inf_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}(\bm{c},c^{\prime})^{n},(\mathcal{D}^{\prime})^{n})\right\}\mid\bm{c}^{\prime}=c^{\prime}\right].$
Using the above and Equation 4, we have that
$\mathop{{\mathds{E}}\/}_{\bm{c}^{\prime}}$ $\left[\inf_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}(\bm{c}^{\prime})^{n},(\mathcal{D}^{\prime})^{n})\right\}\right]$
$\quad\quad\leq\mathop{{\mathds{E}}\/}_{\bm{c}^{\prime}}\left[\mathop{{\mathds{E}}\/}_{\bm{c}\mid\bm{c}^{\prime}}\left[\inf_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{\sqrt{\frac{n\ln(1/p(\bm{c},\bm{c}^{\prime}))])}{2(m-k)}}+\frac{nk}{m}\right\}\right]\right]$
$\quad\quad=\frac{nk}{m}+\sqrt{\frac{n}{2(m-k)}}\cdot\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\sum_{c^{\prime},\text{ where }(\bm{c},c^{\prime})\text{ are feasible}}p(\bm{c},c^{\prime})\sqrt{\ln(1/p(\bm{c},c^{\prime}))}\right].$
For any fixed choice of $c$, there are at most $d^{k}$ feasible $c^{\prime}$. The summation, therefore, contains at most $d^{k}$ terms, and the sum of $p(c,c^{\prime})$ over these $d^{k}$ terms is $1$. Since the function $x\mapsto x\sqrt{\ln(1/x)}$ is concave, this summation is maximized if all the $p(c,c^{\prime})$ are the same value of $d^{-k}$, and so,
$\sum_{c^{\prime},\text{ where }(c,c^{\prime})\text{ are feasible}}p(c,c^{\prime})\sqrt{\ln(1/p(c,c^{\prime}))}\leq d^{k}\cdot d^{-k}\sqrt{\ln(1/d^{-k})}=\sqrt{k\ln d}.$
Therefore, we conclude that,
$\mathop{{\mathds{E}}\/}_{\bm{c}^{\prime}}\left[\inf_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}(\bm{c}^{\prime})^{n},(\mathcal{D}^{\prime})^{n})\right\}\right]\leq\sqrt{\frac{nk\ln d}{2(m-k)}}+\frac{nk}{m}.\qed$
7 Bounding the error due to grouping
In this section, we prove the following, restated for convenience. See 4.4 Formally, the quantity $\mathrm{info}_{r}(\tilde{\mathcal{D}})$ measures the mutual information between one uniform coordinate of $\bm{S}$ and $r$ other uniform coordinates. It is defined as
$\mathrm{info}_{r}(\tilde{\mathcal{D}})\coloneqq I(\bm{\sigma}(\bm{S})_{\leq r};\bm{\sigma}(\bm{S})_{r+1}),$
where $\bm{S}\sim\tilde{\mathcal{D}}$, $\bm{\sigma}:[m]\to[m]$ is a uniform permutation, and $\sigma(S)$ indicates the element of $X^{m}$ satisfying $\sigma(S)_{i}=S_{\sigma(i)}$. To avoid the need to track $\bm{\sigma}$, we will restate Lemma 4.4 for permutation invariant $\tilde{\mathcal{D}}$.
Definition 13 (Permutation invariant).
We say a random variable $\bm{S}$ over $X^{m}$ is permutation invariant if, for any fixed $S\in X^{m}$ and permutation $\sigma:[m]\to[m]$,
$\operatorname{{Pr}}[\bm{S}=S]=\operatorname{{Pr}}[\bm{S}=\sigma(S)].$
Similarly, we say a distribution $\mathcal{D}$ with support $X^{m}$ is permutation invariant if $\bm{S}\sim\mathcal{D}$ is permutation invariant.
Lemma 7.1.
For any permutation invariant distribution $\tilde{\mathcal{D}}$ over $X^{m}$ let $\mathcal{D}_{\mathrm{goal}}(c)$ be the distribution of $\bm{x}\sim\mathrm{Unif}(\bm{P})$ where $\bm{P}\sim\tilde{\mathcal{P}}(c)$. For any integers $n+k_{\max}<m$, there exists some $k\in[0,k_{\max}]$ for which, for $\bm{S}\sim\tilde{\mathcal{D}}$ and $\bm{c}\sim\Phi_{m\to k}(\bm{S})$,
$d_{\mathrm{KL}}\left({\mathop{{\mathds{E}}\/}_{\bm{S}}\left[\Phi_{m\to n}(\bm{S})\right]}\,\middle\|\,{\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}\right]}\right)\leq\frac{n^{2}}{2k_{\max}}\cdot I(\bm{S}_{<n+k_{\max}};\bm{S}_{n+k_{\max}}).$
Comparing Lemma 7.1 to Lemma 4.4, the differences are:
  1. 1.
    Lemma 7.1 gives a bound on the KL-divergence whereas Lemma 4.4 gives a bound on the TV-distance. To get the TV bound, we use Pinkser’s inequality (Fact 5.8).
  2. 2.
    Lemma 7.1 requires $\tilde{\mathcal{D}}$ to be permutation invariant and uses $I(\bm{S}_{\leq r};\bm{S}_{r+1})$ in place of $\mathrm{info}_{r}(\tilde{\mathcal{D}})$. We observe these two quantities are the same for permutation invariant distributions. Therefore, to recover Lemma 4.4 from Lemma 7.1, we apply it with $\tilde{\mathcal{D}}$ the distribution of $\bm{\sigma}(\bm{S})$ where $\bm{\sigma}$ is a uniform permutation and $\bm{S}$ is drawn from the $\tilde{\mathcal{D}}$ specified in Lemma 4.4.
Lemma 7.1 is an easy consequence of the following bound on the desired KL divergence.
Lemma 7.2.
For integers $n+k<m$ and permutation invariant distribution $\tilde{\mathcal{D}}$ over $X^{m}$, let $\mathcal{D}_{\mathrm{goal}}(c)$ be the distribution of $\bm{x}\sim\mathrm{Unif}(\bm{P})$ where $\bm{P}\sim\tilde{\mathcal{P}}(c)$. Then, for $\bm{S}\sim\tilde{\mathcal{D}}$ and $\bm{c}\sim\Phi_{m\to k}(\bm{S})$,
$d_{\mathrm{KL}}\left({\mathop{{\mathds{E}}\/}_{\bm{S}}\left[\Phi_{m\to n}(\bm{S})\right]}\,\middle\|\,{\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}\right]}\right)\leq\sum_{i\in[n-1]}I(\bm{S}_{\leq i};\bm{S}_{[m-k,m]})-I(\bm{S}_{\leq i};\bm{S}_{[m-k+1,m]}).$
Proof of Lemma 7.1 given Lemma 7.2.
Summing up the KL divergences from Lemma 7.2 for all $k=0,\ldots,k_{\max}$, cancelling the telescoping terms, and observing that, for $k=0$, the second term vanishes, we obtain
$\sum_{k\in[0,k_{\max}]}\sum_{i\in[n]}I(\bm{S}_{\leq i};\bm{S}_{[m-k,m]})-I(\bm{S}_{\leq i};\bm{S}_{[m-k+1,m]})\leq\sum_{i\in[n-1]}I(\bm{S}_{\leq i};\bm{S}_{[m-k_{\max},m]}).$
Each of these terms can be decomposed using the chain rule (Fact 5.9),
$I(\bm{S}_{\leq i};\bm{S}_{[m-k_{\max},m]})$ $=\sum_{j=1}^{i}I(\bm{S}_{j};\bm{S}_{[m-k_{\max},m]}\mid\bm{S}_{<j})$
$=\sum_{j=1}^{i}I(\bm{S}_{j};\bm{S}_{[m-k_{\max},m]},\bm{S}_{<j})-I(\bm{S}_{j};\bm{S}_{<j})$ (Fact 5.9 again)
$\leq\sum_{j=1}^{i}I(\bm{S}_{j};\bm{S}_{[m-k_{\max},m]},\bm{S}_{<j})$ (Fact 5.10)
Each term in the above expression is the mutual information between one element of $\bm{S}$ and $k+j$ other elements. By permutation invariance of $\bm{S}$, it doesn’t matter which elements we choose. Therefore, the prior sum is bounded by
$\sum_{k\in[0,k_{\max}]}\sum_{i\in[n]}$ $I(\bm{S}_{\leq i};\bm{S}_{[m-k,m]})-I(\bm{S}_{\leq i};\bm{S}_{[m-k+1,m]})$
$\leq\sum_{i\in[n-1]}\sum_{j=1}^{i}I(\bm{S}_{\leq k+j};\bm{S}_{n+k_{\max}})$
$\leq\sum_{i\in[n-1]}\sum_{j=1}^{i}I(\bm{S}_{\leq n+{k_{\max}-1}};\bm{S}_{n+k_{\max}})$ (Fact 5.11)
$=\frac{n(n-1)}{2}I(\bm{S}_{\leq n+{k_{\max}-1}};\bm{S}_{n+k_{\max}}).$
Therefore, using the fact that minimum over all $k\in[0,k_{\max}]$ is at most the mean, there exists one choice of $k$ for which
$d_{\mathrm{KL}}\left({\mathop{{\mathds{E}}\/}_{\bm{S}}\left[\Phi_{m\to n}(\bm{S})\right]}\,\middle\|\,{\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}\right]}\right)\leq\frac{n(n-1)}{2(k_{\max}+1)}\cdot I(\bm{S}_{<n+k_{\max}};\bm{S}_{n+k_{\max}}).\qed$
The proof of Lemma 7.2 uses the following.
Proposition 7.3.
Let $\mathcal{D}$ be any distribution over $X^{n}$ and $\bm{x}\sim\mathcal{D}$. Then,
$d_{\mathrm{KL}}\left({\mathcal{D}}\,\middle\|\,{\mathcal{D}_{1}\times\mathcal{D}_{2}\times\cdots\mathcal{D}_{n}}\right)=\sum_{i=1}^{n-1}I(\bm{x}_{\leq i};\bm{x}_{i+1}).$
Proof.
Throughout this proof, we use $\mathcal{D}_{i}$ to denote the marginal distribution over the $i^{\text{th}}$ coordinate of $\mathcal{D}$ and $\mathcal{D}_{\leq i}$ the marginal distribution over the first $i$ coordinates of $\mathcal{D}$. Expanding the right-hand side,
$\sum_{i=1}^{n-1}I(\mathcal{D}_{\leq i};\mathcal{D}_{i+1})$ $=\sum_{i=1}^{n-1}d_{\mathrm{KL}}\left({\mathcal{D}_{\leq i+1}}\,\middle\|\,{\mathcal{D}_{\leq i}\times\mathcal{D}_{i+1}}\right)$ (Definition of mutual information)
$=\sum_{i=1}^{n-1}\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{\leq i+1}}\left[\ln\left(\frac{\mathcal{D}_{\leq i+1}(\bm{x})}{\mathcal{D}_{\leq i}(\bm{x}_{\leq i})\mathcal{D}_{i+1}(\bm{x}_{i+1})}\right)\right]$ (Definition of KL divergence)
$=\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}\left[\ln\left(\prod_{i=1}^{n-1}\frac{\mathcal{D}_{\leq i+1}(\bm{x}_{\leq i+1})}{\mathcal{D}_{\leq i}(\bm{x}_{\leq i})\mathcal{D}_{i+1}(\bm{x}_{i+1})}\right)\right]$ (Linearity of expectation)
$=\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}\left[\ln\left(\frac{\mathcal{D}(\bm{x})}{\prod_{i=1}^{n}\mathcal{D}_{i}(\bm{x}_{i})}\right)\right]$ (Cancellation of terms)
which is exactly $d_{\mathrm{KL}}\left({\mathcal{D}}\,\middle\|\,{\mathcal{D}_{1}\times\mathcal{D}_{2}\times\cdots\mathcal{D}_{n}}\right)$. ∎
We are now ready to prove Lemma 7.2.
Proof.
For any $c\in X^{k}$, let $\mathcal{D}_{\mathrm{true}}$ be the distribution of $\Phi_{(m-k)\to n}(\bm{P})$ for $\bm{P}\sim\tilde{\mathcal{P}}(c)$. We first claim that $\mathop{{\mathds{E}}\/}_{\bm{S}}[\Phi_{m\to n}(\bm{S})]$ and $\mathop{{\mathds{E}}\/}_{\bm{c}}[\mathcal{D}_{\mathrm{true}}(\bm{c})]$ are the same distribution.
To prove this claim, suppose we draw $\bm{S}\sim\tilde{\mathcal{D}}$ and $\bm{c}\sim\Phi_{m\to k}(\bm{S})$ and then set $\bm{P}=\bm{S}\setminus\bm{c}$. Then, by Definition 8, this is equivalent to if we had first drawn $\bm{c}$ and then $\bm{P}\sim\tilde{\mathcal{P}}(\bm{c})$. Therefore, we only need to argue that $\Phi_{m\to n}(\bm{S})$ and $\Phi_{m-k\to n}(\bm{P})$ have the same distribution. For this, observe that if we don’t condition on the value of $\bm{c}$, that $\Phi_{m-k\to n}(\bm{P})$ is equally likely to be any $n$ points drawn without replacement from $\bm{S}$, as desired.
Using this equivalence, we have that
$d_{\mathrm{KL}}\left({\mathop{{\mathds{E}}\/}_{\bm{S}}\left[\Phi_{m\to n}(\bm{S})\right]}\,\middle\|\,{\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}\right]}\right)=d_{\mathrm{KL}}\left({\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{true}}(\bm{c})\right]}\,\middle\|\,{\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}\right]}\right).$
Then, using the convexity of KL divergence (Fact 5.7) and Jensen’s inequality,
$d_{\mathrm{KL}}\left({\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{true}}(\bm{c})\right]}\,\middle\|\,{\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}\right]}\right)\leq\mathop{{\mathds{E}}\/}_{\bm{c}}\left[d_{\mathrm{KL}}\left({\mathcal{D}_{\mathrm{true}}(\bm{c})}\,\middle\|\,{\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}}\right)\right].$
We then observe that, for any fixed choice of $c$, the marginal distribution of $\mathcal{D}_{\mathrm{true}}(c)$ on each of the $n$ coordinates is exactly equal to $\mathcal{D}_{\mathrm{goal}}(c)$. We then apply Proposition 7.3, using $\bm{P}$ to represent a permutation invariant draw from $\mathcal{P}(\bm{c})$,
$\mathop{{\mathds{E}}\/}_{\bm{c}}\left[d_{\mathrm{KL}}\left({\mathcal{D}_{\mathrm{true}}(\bm{c})}\,\middle\|\,{\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}}\right)\right]=\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\sum_{i=1}^{n-1}I(\bm{P}_{\leq i};\bm{P}_{i+1}\mid\bm{c})\right].$
Then, use the chain rule (Fact 5.9),
$I(\bm{P}_{\leq i};\bm{P}_{i+1}\mid\bm{c})=I(\bm{P}_{\leq i};\bm{P}_{i+1},\bm{c})-I(\bm{P}_{\leq i};\bm{c}).$
Finally, recall that the distribution of $\bm{P}\circ\bm{c}$ is the same as that of $\bm{S}$. Therefore, using the permutation invariance of $\bm{S}$,
$d_{\mathrm{KL}}\left({\mathop{{\mathds{E}}\/}_{\bm{S}}\left[\Phi_{m\to n}(\bm{S})\right]}\,\middle\|\,{\mathop{{\mathds{E}}\/}_{\bm{c}}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n}\right]}\right)\leq\sum_{i=1}^{n-1}I(\bm{S}_{\leq i};\bm{S}_{[m-k,m]})-I(\bm{S}_{\leq i};\bm{S}_{[m-k+1,m]}).\qed$
7.1 Bounding the mutual information for low-degree cost functions
We prove the following, restated for convenience. See 4.5
Proof.
Throughout this proof, we’ll use ${\bm{T}}$ and ${\bm{T}}^{\prime}$ as shorthand for $\bm{\sigma}(\bm{S})$ and $\bm{\sigma}(\textsc{Corrupt}(\bm{S}))$ for $\bm{\sigma}:[m]\to[m]$ a uniform permutation. Then,
$\mathrm{info}_{r}(\tilde{\mathcal{D}})$ $=I({\bm{T}}^{\prime}_{\leq r};{\bm{T}}_{r+1})$
$\leq I\left(({\bm{T}}_{\leq r},{\bm{T}}_{\leq r}^{\prime});({\bm{T}}_{r+1},{\bm{T}}_{r+1}^{\prime})\right)$ (Fact 5.11)
$=I\left(({\bm{T}}_{\leq r},{\bm{T}}_{\leq r}^{\prime});{\bm{T}}_{r+1}\right)+I\left(({\bm{T}}_{\leq r},{\bm{T}}_{\leq r}^{\prime});{\bm{T}}_{r+1}^{\prime}\mid{\bm{T}}_{r+1}\right)$ (Fact 5.9)
$=I({\bm{T}}_{\leq r},;{\bm{T}}_{r+1})+I({\bm{T}}_{\leq r}^{\prime};{\bm{T}}_{r+1}\mid{\bm{T}}_{\leq r})+I\left(({\bm{T}}_{\leq r},{\bm{T}}_{\leq r}^{\prime});{\bm{T}}_{r+1}^{\prime}\mid{\bm{T}}_{r+1}\right).$ (Fact 5.9)
The first term, $I({\bm{T}}_{\leq r},;{\bm{T}}_{r+1})$, is zero because ${\bm{T}}_{\leq r}$ and ${\bm{T}}_{r+1}$ are independent. The third term, $I\left(({\bm{T}}_{\leq r},{\bm{T}}_{\leq r}^{\prime});{\bm{T}}_{r+1}^{\prime}\mid{\bm{T}}_{r+1}\right)$, is at most $\ln(d)$ by Fact 5.13 and the fact that conditioned on the value of ${\bm{T}}_{r+1}$, there are only $d$ possible values of ${\bm{T}}_{r+1}^{\prime}$. To bound the remaining term, we will show that
$I({\bm{T}}_{\leq r}^{\prime};{\bm{T}}_{r+1}\mid{\bm{T}}_{\leq r})=\frac{1}{m-r}\cdot\sum_{i=r+1}^{n}I({\bm{T}}_{\leq r}^{\prime};{\bm{T}}_{i}\mid{\bm{T}}_{\leq r})\leq\frac{1}{m-r}\cdot I({\bm{T}}_{\leq r}^{\prime};{\bm{T}}_{>r}\mid{\bm{T}}_{\leq r})\leq\frac{r\ln d}{m-r}.$
In the above, the equality holds by permutation invariance, and the second inequality follows from Fact 5.13 and that, conditioned on the value of ${\bm{T}}_{\leq r}$, there are only $d^{r}$ choices for ${\bm{T}}_{\leq r}^{\prime}$. For the first inequality,
$I({\bm{T}}_{\leq r}^{\prime}$ $;{\bm{T}}_{>r}\mid{\bm{T}}_{\leq r})$
$=\sum_{i=r+1}^{n}I\left({\bm{T}}_{\leq r}^{\prime};{\bm{T}}_{i}\mid{\bm{T}}_{\leq r},{\bm{T}}_{[r+1,i-1]}\right)$ (Chain rule, Fact 5.9)
$=\sum_{i=r+1}^{n}I\left(({\bm{T}}_{\leq r}^{\prime},{\bm{T}}_{[r+1,i-1]});{\bm{T}}_{i}\mid{\bm{T}}_{\leq r}\right)-I\left({\bm{T}}_{[r+1,i-1]};{\bm{T}}_{i}\mid{\bm{T}}_{\leq r}\right)$ (Chain rule again)
$=\sum_{i=r+1}^{n}I\left(({\bm{T}}_{\leq r}^{\prime},{\bm{T}}_{[r+1,i-1]});{\bm{T}}_{i}\mid{\bm{T}}_{\leq r}\right)$ (All coordinates of ${\bm{T}}$ are independent)
$\geq\sum_{i=r+1}^{n}I\left({\bm{T}}_{\leq r}^{\prime};{\bm{T}}_{i}\mid{\bm{T}}_{\leq r}\right).$ (Fact 5.11)
Combining all the above, we have that,
$\mathrm{info}_{r}(\tilde{\mathcal{D}})\leq\ln d+\frac{r\ln d}{m-r}=\frac{m\ln d}{m-r}.\qed$
8 Adaptive adversaries are at least as strong as oblivious adversaries
In this version, we prove the easy direction of Theorem 6. Such a statement is well-known to hold for some specific [DKK+19, ZJS19] adversary models. Here, we show it holds with any cost function using a simulation argument. See 4.1 The high-level idea in the proof of Proposition 4.1 is simple: Since $\mathcal{D}^{\prime}$ is a valid corruption of $\mathcal{D}$, it is possible to couple samples $\bm{S}\sim\mathcal{D}^{n}$ and $\bm{S}^{\prime}\sim(\mathcal{D}^{\prime})$ so that the average cost of corrupting a point in $\bm{S}_{i}$ to the corresponding point in $\bm{S}^{\prime}_{i}$ is at most $1$. If, the adaptive adversary could set $\textsc{Corrupt}(\bm{S})$ to the value of random variable $\bm{S}^{\prime}\mid\bm{S}$, then it would easily simulate this oblivious adversary. However, this may not be a valid corruption because, even though the corruption of $\bm{S}$ to $\bm{S}^{\prime}$ has an average cost of $1$, it can exceed $1$.
Therefore, our strategy will be to round $\bm{S}^{\prime}$ to a valid corruption. The quantity we need to bound is how many points of $\bm{S}^{\prime}$ do we need to change to make the corruption valid, which we do in the following lemma.
Claim 8.1.
For any distribution $\mathcal{C}$ supported on $R_{\geq 0}$ with mean $1$, draw $\bm{x}_{1},\ldots,\bm{x}_{n}\overset{\mathrm{iid}}{\sim}\mathcal{C}$ and define $\Delta(x_{1},\ldots,x_{n})$ to be the minimum number of $x_{i}$ that must be removed so that the sum of the remaining elements is at most $n$. Then,
$\mathop{{\mathds{E}}\/}[\Delta(\bm{x}_{1},\ldots,\bm{x}_{n})]\leq 5\sqrt{n}.$
The main ingredient in the proof of Claim 8.1 is an upper bound on the probability that $\Delta(\bm{x}_{1},\ldots,\bm{x}_{n})$ exceeds a value.
Proposition 8.2.
In the setting of Claim 8.1, for any $v\geq 0$
$\operatorname{{Pr}}\left[\Delta(\bm{x}_{1},\ldots,\bm{x}_{n})\geq v\right]\leq\frac{2}{v}+\frac{4n}{v^{2}}.$
Proof.
The main idea in this proof is, for any $r\in[0,1]$, to exhibit a strategy with the following properties.
  1. 1.
    The probability the strategy removes more than $2nr$ elements is at most $\frac{1}{nr}$.
  2. 2.
    The probability that, after this strategy removes elements, the remaining sum is more than $n$ is at most $\frac{1}{nr^{2}}$.
Combining the above with a union bound gives that
$\operatorname{{Pr}}[\Delta(\bm{x})\geq 2nr]\leq\frac{1}{nr}+\frac{1}{nr^{2}}.$
The above is equivalent to the desired result as we can set $r=\frac{v}{2n}$.
For any $\tau\in\mathds{R}$, $p\in[0,1]$ consider the strategy that, for each $i\in[m]$ keeps $\bm{x}_{i}$ with probability $f(\bm{x}_{i})$ and otherwise removes it for
$f(x)\coloneqq\begin{cases}1&\text{if }x<\tau\\ p&\text{if }x=\tau\\ 0&\text{if }x>\tau.\end{cases}$
It is always possible to choose $\tau$ and $p$ so that the probability $\bm{x}_{i}$ is removed is any desired value. We set them so that the probability $\bm{x}_{i}$ is removed is exactly $r$.
It is unlikely many elements are removed. Let ${\bm{R}}$ be the random variable representing the number of removed elements. Then, since the probability each element is removed is $r$,
$\mathop{{\mathds{E}}\/}[{\bm{R}}]=nr.$
Furthermore, since the elements are sampled without replacement, the indicators of whether each element is removed are negatively correlated. Therefore,
$\operatorname{{Var}}[{\bm{R}}]\leq n\operatorname{{Var}}[\mathds{1}[\bm{x}_{i}\text{ removed}]]=nr(1-r)\leq nr.$
Lastly, we apply Chebyshev’s inequality:
$\operatorname{{Pr}}[{\bm{R}}\geq 2nr]\leq\frac{\operatorname{{Var}}[{\bm{R}}]}{(2nr-\mathop{{\mathds{E}}\/}[{\bm{R}}])^{2}}\leq\frac{nr}{(nr)^{2}}=\frac{1}{nr}.$
It is unlikely the sum of the remaining elements is more than $n$. Let $\bm{X}$ be the sum of the remaining elements. Then,
$\bm{x}\coloneqq\sum_{i\in[n]}\bm{z}_{i}\cdot\bm{x}_{i}\quad\quad\text{for}\quad\bm{z}_{i}\sim\mathrm{Ber}(f(\bm{x}_{i})).$
We will analyze the mean and variance of this $\bm{X}$ and then use Chebyshev’s inequality to bound the probability it is more than $1$. For the mean,
$\mathop{{\mathds{E}}\/}[\bm{X}]$ $=n\cdot\mathop{{\mathds{E}}\/}[\bm{z}_{i}\cdot\bm{x}_{i}]$ (Linearity of expectation)
$=n\cdot\left(\mathop{{\mathds{E}}\/}[\bm{x}_{i}]-\operatorname{{Pr}}[\bm{z}_{i}=0]\cdot\mathop{{\mathds{E}}\/}[\bm{x}_{i}\mid\bm{z}_{i}=0]\right)$ ($\bm{z}_{i}$ supported on $\{0,1\}$)
$\leq n\cdot\left(1-r\cdot\tau\right).$
For the variance of $\bm{X}$, since $\bm{x}_{1},\ldots,\bm{x}_{m}$ are independent, the variances sum. Therefore,
$\operatorname{{Var}}[\bm{X}]$ $=n\cdot\operatorname{{Var}}[\bm{z}_{i}\cdot\bm{x}_{i}]$
$\leq n\cdot\mathop{{\mathds{E}}\/}[(\bm{z}_{i}\cdot\bm{x}_{i})\cdot(\bm{z}_{i}\cdot\bm{x}_{i})]$
$\leq n\cdot\max(\bm{z}_{i}\cdot\bm{x}_{i})\cdot\mathop{{\mathds{E}}\/}[\bm{z}_{i}\cdot\bm{x}_{i}]$
$\leq n\tau.$
Therefore, by applying Chebyshev’s inequality, we see that
$\operatorname{{Pr}}[\bm{X}\geq n]\leq\frac{n\tau}{n^{2}\tau^{2}r^{2}}=\frac{1}{n\tau^{2}r^{2}}.$
Note furthermore that if $\tau<1$, then every term in the sum is less than $1$, in which case $\operatorname{{Pr}}[\bm{P}\geq 1]=0$. Therefore, the worst-case choice of $\tau$ for our bound is $\tau=1$ in which case
$\operatorname{{Pr}}[\bm{P}>n]\leq\frac{1}{nr^{2}}.\qed$
Proof of Claim 8.1.
We write,
[…]
$=2\sqrt{n}+2\sqrt{n}+\sqrt{n}=5\sqrt{n}.\qed$
We are now ready to prove the main result of this section.
Proof of Proposition 4.1.
Consider any $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$. By definition, there is a coupling between $\bm{x}\sim\mathcal{D}$ and $\bm{y}\sim\mathcal{D}^{\prime}$ so that $\mathop{{\mathds{E}}\/}[\rho(\bm{x},\bm{y})]\leq 1$. We can extend this to a coupling of $\bm{S}\sim\mathcal{D}^{m}$ and $\bm{S}^{\prime}\sim(\mathcal{D}^{\prime})^{m}$ so that
$\mathop{{\mathds{E}}\/}\left[\frac{1}{m}\cdot\sum_{i\in m]}\rho(\bm{S}_{i},\bm{S}^{\prime}_{i})\right]\leq 1.$
Let $\mathcal{C}$ be the distribution of $\rho(\bm{x},\bm{y})$ in this coupling. By Claim 8.1, we can construct $\bm{S}^{\prime\prime}$ for which $\frac{1}{m}\cdot\sum_{i\in m]}\rho(\bm{S}_{i},\bm{S}^{\prime\prime}_{i})\leq 1$ with probability $1$ and for which the expected number of coordinates on which $\bm{S}^{\prime\prime}$ and $\bm{S}^{\prime}$ differ is at most $5\sqrt{m}$. This is because, whenever Claim 8.1 asks to “remove” some $\bm{x}_{i}$, we can simply set $\bm{S}_{i}^{\prime\prime}=\bm{S}_{i}$ in which case $\rho(\bm{S}_{i},\bm{S}_{i}^{\prime\prime})=0$. We set Corrupt to the (possibly randomized) [^3] In some sense, this adaptive adversary only needs to be randomized if the oblivious adversary is randomized. Specifically, we could define a deterministic oblivious adversary as one choosing a deterministic function $\textsc{Corrupt}:X\to X$ satisfying $\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}[\rho(\bm{x},\textsc{Corrupt}(\bm{x}))]\leq 1$ and which corresponds to $\mathcal{D}^{\prime}$ being the distribution of $\textsc{Corrupt}(\bm{x})$. For such an adversary, the values of $\bm{S}^{\prime}$ and $\bm{S}^{\prime\prime}$ will be deterministic functions of $\bm{S}$, and so the adaptive adversary is also determinstic. function mapping $\bm{S}$ to $\bm{S}^{\prime\prime}\mid\bm{S}$. Then, for any $f:X^{n}\to[0,1]$
$\mathop{{\mathds{E}}\/}[f\circ\Phi_{{m}\to n}(\textsc{Corrupt}(\bm{S}))]$ $=\mathop{{\mathds{E}}\/}[f\circ\Phi_{{m}\to n}(\bm{S}^{\prime\prime})]$
$\geq\mathop{{\mathds{E}}\/}[f\circ\Phi_{{m}\to n}(\bm{S}^{\prime})]-\operatorname{{Pr}}[\text{$\bm{S}^{\prime\prime}$ differs from $\bm{S}^{\prime}$ on one of $n$ sampled points}]$
$\geq\mathop{{\mathds{E}}\/}[f\circ\Phi_{{m}\to n}(\bm{S}^{\prime})]-\frac{5n}{\sqrt{m}}.$
The desired result follows from our choice of $m$ and that the distribution of $\Phi_{{m}\to n}(\bm{S}^{\prime})$ for $\bm{S}^{\prime}\sim(\mathcal{D}^{\prime})^{m}$ is simply $(\mathcal{D}^{\prime})^{n}$. ∎
9 Randomized simulations and indistinguishability are equivalent
We prove the following, restated for convenience. See 4.2 We prove each direction separately. The first direction is nearly immediate.
Proposition 9.1 (Randomized simulations imply indistinguishability).
In the setting of Lemma 4.2, randomized simulations imply indistinguishability even when $X$ and/or $\mathscr{D}_{2}$ are not finite.
Proof.
Consider any $\mathcal{D}_{1}\in\mathscr{D}_{1}$ and $f:X\to[0,1]$. We wish to show there is some $\mathcal{D}_{2}\in\mathscr{D}_{2}$ for which
$\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{1}}[f(\bm{x})]\leq\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{2}}[f(\bm{x})]+\varepsilon.$
By the randomized simulation property, there exists $\bm{\mathcal{D}}_{2}$ supported on $\mathscr{D}_{2}$ for which
$d_{\mathrm{TV}}(\mathcal{D}_{1},\mathop{{\mathds{E}}\/}[\bm{\mathcal{D}}_{2}])\leq\varepsilon.$
By the definition of total variation distance,
$\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{1}}[f(\bm{x})]-\mathop{{\mathds{E}}\/}_{\bm{\mathcal{D}}_{2}}\left[\mathop{{\mathds{E}}\/}_{\bm{x}\sim\bm{\mathcal{D}}_{2}}[f(\bm{x})]\right]\leq\varepsilon.$
The desired result follows from the fact that there must exist some $\mathcal{D}_{2}$ in the support of $\bm{\mathcal{D}}_{2}$ for which $\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}_{2}}[f(\bm{x})]$ is at least its mean value over the choice of $\bm{\mathcal{D}}_{2}$. ∎
For the other direction, we will apply Von Neumann’s minimax theorem.
Fact 9.2 (The minimax theorem [vN28]).
Let $L(x,y)$ be a bilinear function (i.e. linear in both $x$ and $y$ when the other is fixed), and $X,Y$ be compact convex sets. Then,
$\max_{x\in X}\min_{y\in Y}L(x,y)=\min_{y\in Y}\max_{x\in X}L(x,y).$
Proposition 9.3.
In the setting of Lemma 4.2, indistinguishability implies randomized simulations when the domain $X$ is finite.
Proof.
Fix any $\mathcal{D}_{1}\in\mathscr{D}_{1}$. For any $f:X\to[0,1]$ and $w\in[0,1]^{\mathscr{D}_{2}}$ let
$L(f,w)=\sum_{x\in X}f(x)\cdot\left(\mathcal{D}_{1}(x)-\sum_{\mathcal{D}_{2}\in\mathscr{D}_{2}}w_{\mathcal{D}_{2}}\cdot\mathcal{D}_{2}(x)\right).$
where $\mathcal{D}(x)$ represents the probability mass function of $\mathcal{D}$ evaluated at $x$. Clearly $L$ is bilinear in $f$ and $w$. Let $F$ be the space of all functions form $X$ to $[0,1]$, and let $W$ be subset of $[0,1]^{\mathscr{D}_{2}}$ that sums to $1$, both of which are compact. Then, by $\varepsilon$-indistinguishability we have that,
$\max_{f\in F}\min_{w\in W}L(f,w)\leq\varepsilon.$
We can therefore apply the minimax theorem to conclude that
$\min_{w\in W}\max_{f\in F}L(f,w)\leq\varepsilon.$
Let $\overline{\mathcal{D}_{2}}$ be the mixture distribution which puts weight $w_{\mathcal{D}_{2}}$ on $\mathcal{D}_{2}$. The above statement says, for any $f:X\to[0,1]$,
$\sum_{x\in X}f(x)\cdot\left(\mathcal{D}_{1}(x)-\overline{\mathcal{D}_{2}}(x)\right)\leq\varepsilon.$
Furthermore, $\overline{\mathcal{D}_{2}}$ is a mixture of distributions from within $\mathscr{D}_{2}$, which is exactly what we wished to show. ∎
10 Putting the pieces together: Proof of Theorem 3
In this section, we combine all the previous ingredients to prove the below theorem, restated for convenience, which also easily implies Theorems 2 and 3. See 6
Proof.
We first prove the easy direction. Consider any corruption by the oblivious adversary, $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$, and let $\textsc{Corrupt}:X^{m}\to X^{m}$ be the function guaranteed to exist by Proposition 4.1. Then,
$\mathop{{\mathds{E}}\/}_{\bm{S}^{\prime}\sim(\mathcal{D}^{\prime})^{n}}[f(\bm{S}^{\prime})]\leq\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{m}}[f\circ\Phi_{{m}\to n}(\textsc{Corrupt}(\bm{S}))]+\varepsilon\leq\mathrm{Adaptive\text{-}Max}_{\rho}(f\circ\Phi_{{m}\to n},\mathcal{D})+\varepsilon,$
where the first inequality is by Proposition 4.1 combined with $m\geq 25n^{2}/\varepsilon^{2}$, and the second is by Corrupt being a valid strategy for the adaptive adversary. The desired result follows by taking a supremum over all $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$.
For the harder direction, we wish to show that for any $f:X^{n}\to\{0,1\}$,
$\mathrm{Adaptive\text{-}Max}_{\rho}(f\circ\Phi_{{m}\to n},\mathcal{D})\leq\mathrm{Oblivious\text{-}Max}_{\rho}(f,\mathcal{D})+\varepsilon.$
By Lemma 4.2, it suffices to show that for any $\textsc{Corrupt}:X^{m}\to X^{m}$ that maps any $S\in X^{m}$ to some $S^{\prime}\in X^{m}$, there is a mixture of oblivious corruptions, $\bm{\mathcal{D}}^{\prime}$, that all fall within $\mathcal{C}_{\rho}(\mathcal{D})$, for which,
$d_{\mathrm{TV}}\left(\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{n}}[\Phi_{{m}\to n}(\textsc{Corrupt}(\bm{S}))],\mathop{{\mathds{E}}\/}_{\bm{\mathcal{D}}^{\prime}}[(\bm{\mathcal{D}}^{\prime})^{n}]\right)\leq\varepsilon.$ (5)
This $\bm{\mathcal{D}}^{\prime}$ will be as described in Figure 1 with whatever $k\leq k_{\max}\coloneqq\lceil\frac{2n^{2}\ln d}{\varepsilon^{2}}\rceil$ minimizes the above mean TV distance. Clearly, every distribution in the support of $\bm{\mathcal{D}}^{\prime}$ is a valid oblivious corruption, because in the last step, we round $\bm{\mathcal{D}}^{\prime}$ to the nearest valid corruption. Therefore, all that remains is the bound this mean TV distance.
For this choice of $k$, we can apply Lemma 4.4 to obtain,
$d_{\mathrm{TV}}$ $\left(\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{m}}\left[\Phi_{{m}\to n}(\textsc{Corrupt}(\bm{S}))\right],\mathop{{\mathds{E}}\/}_{\bm{S},\bm{c}\sim\Phi_{m\to k}(\textsc{Corrupt}(\bm{S}))}\left[\mathcal{D}_{\mathrm{goal}}(\bm{c})\right]\right)$
$\quad\quad\leq\sqrt{\frac{n^{2}\cdot\mathrm{info}_{n+k-1}(\tilde{\mathcal{D}})}{4k_{\max}}}$ (Lemma 4.4)
$\quad\quad\leq\sqrt{\frac{2n^{2}\ln d}{4k_{\max}}}$ (Lemma 4.5 and $k+n\leq m/2$)
$\quad\quad\leq\frac{\varepsilon}{2}.$ (our choice of $k_{\max}\coloneqq\lceil\frac{2n^{2}\ln d}{\varepsilon^{2}}\rceil$)
Next, by Lemma 6.5,
$\mathop{{\mathds{E}}\/}_{\bm{c}\sim\Phi_{m\to k}(\textsc{Corrupt}(\bm{S}))}\left[\inf_{\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})}\left\{d_{\mathrm{TV}}(\mathcal{D}_{\mathrm{goal}}(\bm{c})^{n},(\mathcal{D}^{\prime})^{n})\right\}\right]\leq\sqrt{\frac{nk\ln d}{2(m-k)}}+\frac{nk}{m}\leq\varepsilon/2,$
where the second inequality uses our choice of $m=k_{\max}(1+\frac{8n\ln d}{\varepsilon^{2}})$ and $k\leq k_{\max}$. Therefore, by the triangle inequality for TV distance, Equation 5 holds. The desired result then follows from Lemma 4.2. ∎
11 Lower bounds
11.1 Proof of Theorem 4
Here, we prove Theorem 4. Given a distribution $\mathcal{D}$ promised to be uniform on some $X^{\prime}\subseteq X\coloneqq[n]$, we show that approximating the cardinality of $X^{\prime}$ is harder in the presence of adaptive subtractive contamination than it is in the presence of oblivious subtractive contamination. To begin, we formalize both models. For ease of notation, we stick with the setting where the adversary can remove half of the sample or distribution, though all the conclusions would remain the same if this $1/2$ were replaced with any other constant.
Definition 14 ($1/2$-Subtractive contamination, special case of Definition 15).
For any distribution $\mathcal{D}$, we say that $\mathcal{D}^{\prime}\in\mathrm{sub}(\mathcal{D})$ if a sample of $\bm{x}^{\prime}\sim\mathcal{D}^{\prime}$ is equivalent to a sample $\bm{x}\sim\mathcal{D}$ conditioned on an event that occurs with probability $1/2$.
Similarly, for any sample $S\in X^{2m}$, we say that $S^{\prime}\in\mathrm{sub}(S)$ if, for $m$ unique indices $i_{1},\ldots,i_{m}\in[2m]$,
$(S^{\prime})_{j}=S_{i_{j}}\quad\quad\text{for all }j\in[m].$
We prove the following.
Theorem 8 (A polynomial increase in sample size is necessary, formal version Theorem 4).
Let $\mathcal{D}$ be a distribution on $X=[n]\coloneqq\{1,\ldots,n\}$ that is promised to be uniform on some $X^{\prime}\subseteq X$. Then for some absolute constant $c<1$,
  1. 1.
    For $m_{\mathrm{small}}\coloneqq O(\sqrt{n})$ and any $k\leq n$, there is an algorithm $f_{\mathrm{obl}}:X^{m_{\mathrm{small}}}\to\{0,1\}$ that distinguishes between the cases where $|X^{\prime}|\geq k$ vs $|X^{\prime}|\leq ck$ with high probability even in the presence of the oblivious adversary,
    $|X^{\prime}|\geq k\implies\inf_{\mathcal{D}^{\prime}\in\mathrm{sub}(\mathcal{D})}\left(\mathop{{\operatorname{{Pr}}}\/}_{\bm{S}\sim(\mathcal{D}^{\prime})^{m_{\mathrm{small}}}}\left[f_{\mathrm{\mathrm{obl}_{\eta}}}(\bm{S})\right]\right)\geq 0.99$
    $|X^{\prime}|\leq ck\implies\sup_{\mathcal{D}^{\prime}\in\mathrm{sub}(\mathcal{D})}\left(\mathop{{\operatorname{{Pr}}}\/}_{\bm{S}\sim(\mathcal{D}^{\prime})^{m_{\mathrm{small}}}}\left[f_{\mathrm{\mathrm{obl}_{\eta}}}(\bm{S})\right]\right)\leq 0.01.$
  2. 2.
    For $m_{\mathrm{large}}=\Omega(n)$ and any $k=n$, there is no algorithm with the same guarantees: Formally, for any $f_{\mathrm{ada}}:X^{m_{\mathrm{large}}}\to\{0,1\}$, either there is an $X^{\prime}$ containing $k$ elements for which
    $\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{2m_{\mathrm{large}}}}\left[\inf_{\bm{S}^{\prime}\in\mathrm{sub}(\bm{S})}\mathds{1}\left[f_{\mathrm{ada}}(\bm{S}^{\prime})\right]\right]\leq 0.51$
    or there is an $X^{\prime}$ containing at most $ck$ elements for which
    $\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{2m_{\mathrm{large}}}}\left[\sup_{\bm{S}^{\prime}\in\mathrm{sub}(\bm{S})}\mathds{1}\left[f_{\mathrm{ada}}(\bm{S}^{\prime})\right]\right]\leq 0.49$
Note that by standard techniques, the above can be converted to a separation in the search version of the problem, where the goal is to approximate $|X^{\prime}|$ to multiplicative accuracy, at the cost of a $\mathrm{polylog}(n)$ dependence.
The construction of $f_{\mathrm{obl}}$ follows from standard results on the probability of a collision in a sample (see e.g. [GR11]).
Fact 11.1 (The probability of a collision in a sample).
Let $\mathcal{D}^{\prime}$ be any distribution supported on $k$ points for which $\mathop{{\operatorname{{Pr}}}\/}_{\bm{x}\sim\mathcal{D}}[\bm{x}=x]\leq\tfrac{2}{k}$ for all possible $x$. Then, for $\bm{x}_{1},\ldots,\bm{x}_{m}\overset{\mathrm{iid}}{\sim}\mathcal{D}^{\prime}$, let $\bm{E}$ indicate whether there is $i\neq j$ for which $\bm{x}_{i}=\bm{x}_{j}$. There are absolute constants $c_{1},c_{2}$ for which,
$c_{1}m^{2}\leq k\implies\operatorname{{Pr}}[\bm{E}]\leq 0.01\quad\quad\text{and}\quad\quad c_{2}m^{2}\geq k\implies\operatorname{{Pr}}[\bm{E}]\geq 0.01.$
The lower bound against the adaptive adversary is an easy consequence of the following simple proposition. If the adversary removes all duplicates, by symmetry, the algorithm cannot distinguish between the case where $|X^{\prime}|$ is large vs small.
Proposition 11.2.
Let $\mathcal{D}$ be uniform on a distribution supported on $n$ points and
$m\coloneqq\lfloor\tfrac{\varepsilon n}{2}\rfloor.$
For $\bm{x}_{1},\ldots,\bm{x}_{2m}\overset{\mathrm{iid}}{\sim}\mathcal{D}$, with probability at least $1-\varepsilon$, the number $i\in[2m]$ for which there exists $j\neq i$ satisfying $\bm{x}_{i}=\bm{x}_{j}$ is at most $m$.
Proof.
Let $\bm{z}_{i}$ be the indicator that there is some $j\neq i$ for which $\bm{x}_{j}=\bm{x}_{i}$. By union bound,
$\mathop{{\mathds{E}}\/}[\bm{z}_{i}]\leq\frac{2m-1}{n}\leq\frac{2m}{n}.$
Applying Markov’s inequality to $\bm{Z}=\sum_{i\in[2m]}\bm{z}_{i}$,
$\operatorname{{Pr}}[\bm{Z}\geq m]\leq\frac{\mathop{{\mathds{E}}\/}[\bm{Z}]}{m}\leq\frac{\frac{2m^{2}}{n}}{m}\leq\frac{2m}{n},$
which is at most $\varepsilon$ for our choice of $m$. ∎
11.2 Proof of Theorem 5
Theorem 9 (Dependence on degree is necessary, formal version of Theorem 5).
For any $b,\delta>0$, large enough $n\in\mathds{N}$ (as a function of $b$ and $\delta$), and cost function $\rho:X\times X\to\mathds{R}_{\geq 0}\cup\{\infty\}$ for which $\rho(x,y)\geq 1+\delta$ whenever $x\neq y$, let
$m=\Omega_{b,\delta}\left(\frac{n}{(\ln n)^{2}}\cdot\ln\deg_{b}(\rho)\right).$
If $m>n$, there is a function $f:X^{n}\to\{0,1\}$ and distribution $\mathcal{D}$ over $X$ for which
$\mathrm{Adaptive\text{-}Max}_{\rho}(f\circ\Phi_{{m}\to n},\mathcal{D})\geq 1-O(1/n)\quad\quad\text{and}\quad\quad\mathrm{Oblivious\text{-}Max}_{\rho}(f,\mathcal{D})\leq O(1/n).$ (6)
We use similar ideas as [BLMT22, Theorem 8], which shows that Theorem 9 holds in the specific case where $\rho$ corresponds to additive noise, though we do need to generalize those ideas to this more general setting.
Let $d$ be the largest integer such that $2^{d}\leq\deg_{b}(\rho)$. By Definition 7, we can choose a subset $X^{\prime}\subseteq X$ of cardinality $2^{d}$ and point $x^{\star}\in X^{\prime}$ for which $\rho(x^{\star},y)\leq b$ for all $y\in X^{\prime}$. Let $M:X\to\{\pm 1\}^{d}\cup\{\vec{0}\}$ be any mapping that takes every element of $X^{\prime}$ to a unique element of {±$\{\pm 1\}^{d}$ and all other elements to $\vec{0}$. For an appropriate threshold $\tau>0$, we’ll define
$f(x_{1},\ldots,x_{n})\coloneqq\begin{cases}1&\text{if for every $x_{i}$, thereis an $x_{j}$ with $j\neq i$ s.t. $\langle M(x_{i}),M(x_{j})\rangle\geq\tau$}\\ 0&\text{otherwise.}\end{cases}$ (7)
This choice of $f$ was analyzed by [BLMT22].
Fact 11.3 (Choosing the threshold $\tau$, [BLMT22]).
For any $p\in(0,1)$ and
$m\geq\Omega_{p}\left(\frac{nd}{(\ln n)^{2}}\right),$
there is a choice of threshold $\tau$ for which both of the following hold:
  1. 1.
    Lemma 7.1 of [BLMT22], a uniform point is hard to correlate with: For any $x_{1},\ldots,x_{n-1}\in X$,
    $\mathop{{\operatorname{{Pr}}}\/}_{\bm{u}\sim\mathrm{Unif}(X^{\prime})}\left[\text{There is an }i\in[n-1]\text{ for which }\langle\bm{u},x_{i}\rangle\geq\tau\right]\leq\frac{1}{n}.$
  2. 2.
    Lemma 7.2 of [BLMT22], an adaptive adversary make all points correlated: Take any $m_{\mathrm{small}}\in\mathds{N}$ for which $pm\leq m_{\mathrm{small}}\leq m$. For $\bm{S}\sim\mathrm{Unif}(X^{\prime})^{m_{\mathrm{small}}}$, there is a strategy for adding $m-m_{\mathrm{small}}$ points to $\bm{S}$ to form $\bm{S}^{\prime}$ for which
    $\mathop{{\mathds{E}}\/}[f\circ\Phi_{{m}\to n}(\bm{S}^{\prime})]\geq 1-\frac{1}{n}.$
Proof of Theorem 9.
Define
$c\coloneqq\min\left(\tfrac{1}{b},1-\tfrac{1}{1+\delta}\right),$
and set $\mathcal{D}$ to the distribution that is equal to $x^{\star}$ with probability $\frac{c}{2}$ and otherwise uniform over $X^{\prime}$,
$\mathcal{D}\coloneqq\tfrac{c}{2}\cdot\{x^{\star}\}+\left(1-\tfrac{c}{2}\right)\cdot\mathrm{Unif}(X^{\prime}).$ (8)
Also, set
$p\coloneqq 1-\frac{c}{4}$
and let $\tau$ be the threshold in Fact 11.3. We will show that both Theorem 9 holds with this choice of $\tau$, $f$ as in Equation 7, and $\mathcal{D}$ as in Equation 8.
We begin by analyzing the oblivious adversary. First, for any $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$, since $\rho(x,x^{\prime})\geq 1+\delta$ for each $x\neq x^{\prime}$, there must be a coupling of $\bm{x}\sim\mathcal{D}$ and $\bm{x}^{\prime}\sim\mathcal{D}^{\prime}$ for which
$\operatorname{{Pr}}[\bm{x}\neq\bm{x}^{\prime}]\leq\frac{1}{1+\delta}$
Furthermore, based on Equation 8, there is a coupling of $\bm{x}\sim\mathcal{D}$ and $\bm{u}\sim\mathrm{Unif}(X^{\prime})$ for which
$\operatorname{{Pr}}[\bm{x}\neq\bm{u}]\leq\frac{c}{2}.$
Combining the above, there is a coupling of $\bm{x}^{\prime}\sim\mathcal{D}^{\prime}$ and $\bm{u}\sim\mathrm{Unif}(X^{\prime})$ for which
$\operatorname{{Pr}}[\bm{x}^{\prime}\neq\bm{u}]\leq\frac{c}{2}+\frac{1}{1+\delta}.$
In particular,
$\operatorname{{Pr}}[\bm{x}^{\prime}=\bm{u}]\geq\left(1-\tfrac{1}{1+\delta}\right)-\frac{c}{2}\geq\frac{c}{2}.$
This means there is a coupling of $\bm{S}\sim\mathcal{D}^{\prime}$ and $\bm{U}\sim\mathrm{Unif}(X^{\prime})$ for which, independently for each $i\in[n]$, with probability at least $c/2$, $\bm{S}_{i}=\bm{U}_{i}$. Because we assumed $n$ is sufficiently large as a function of $b$ and $\delta$, we are free to assume that $c\geq\Omega((\ln n)/n)$. As a result, with probability at least $1-1/n$, there is some $i\in[n]$ for which $\bm{S}_{i}=\bm{U}_{i}$. Then,
$\mathop{{\mathds{E}}\/}_{\bm{S}\sim(\mathcal{D}^{\prime})^{n}}[f(\bm{S})]\leq\frac{1}{n}+\mathop{{\mathds{E}}\/}_{\bm{S}}[f(\bm{S})\mid\bm{S}_{i}=\bm{U}_{i}\text{ for some }i\in[n]]\leq\frac{2}{n},$
where the second inequality is by the first part of Fact 11.3.
We proceed to analyze the adaptive adversary. To draw $\bm{S}\sim\mathcal{D}^{m}$, we can first draw $\bm{E}\sim\mathrm{Ber}(c/2)^{m}$. Then, for each $i\in[m]$, if $\bm{E}_{i}=1$ we set $\bm{S}=x^{\star}$ and otherwise draw it uniformly from $X^{\prime}$.
Conditioned on $\sum_{i}\bm{E}_{i}=E$, we have that $E$ of the elements in $\bm{S}$ are set to $x^{\star}$ and the other $m-E$ are drawn independently and uniformly from $X^{\prime}$. By the second part of Fact 11.3, whenever $m-E\geq pm$ (or equivalently, $E\leq\frac{mc}{4}$), there is a way to modify the $E$ many elements for which $\bm{E}_{i}=1$ to form a corrupted sample $\bm{S}^{\prime}$ satisfying
$\mathop{{\mathds{E}}\/}[f\circ\Phi_{{m}\to n}(\bm{S}^{\prime})]\geq 1-\frac{1}{n}.$
Furthermore, if $E\leq mc\leq\frac{m}{b}$, the adversary has enough budget to modify all of the indices for which $\bm{E}_{i}=1$ to arbitrary elements of $X^{\prime}$. Therefore, there is a strategy for the adaptive adversary so that
$\mathop{{\mathds{E}}\/}\left[f\circ\Phi_{{m}\to n}(\bm{S}^{\prime})\,\bigg{|}\,\frac{mc}{4}\leq\sum_{i\in[m]}\bm{E}_{i}\leq mc\right]\geq 1-\frac{1}{n}.$
We once again assume that $c\geq\Omega((\ln n)/n)$, which implies that $c\geq\Omega((\ln m)/m)$. By a Chernoff bound, this gives that
$\operatorname{{Pr}}\left[\tfrac{mc}{4}\leq\sum_{i\in[m]}\bm{E}_{i}\leq mc\right]\geq 1-\frac{1}{n}.$
So by union bound,
$\mathop{{\mathds{E}}\/}\left[f\circ\Phi_{{m}\to n}(\bm{S}^{\prime})\right]\geq 1-\frac{2}{n}.\qed$
12 Acknowledgments
The authors thank Li-Yang Tan for his helpful discussions and feedback. Gregory is supported by a Simons Foundation Investigator Award, NSF award AF-2341890 and UT Austin’s Foundations of ML NSF AI Institute. Guy is supported by NSF awards 1942123, 2211237, and 2224246 and a Jane Street Graduate Research Fellowship.
Appendix A The subtractive and additive adversaries in our framework
We show how to fit subtractive and additive contamination into our framework.
A.1 Subtractive adversaries
First, we formally define $\eta$-subtractive corruptions
Definition 15.
For any distribution $\mathcal{D}$ and $\eta\in(0,1)$, we say that $\mathcal{D}^{\prime}$ is an $\eta$-subtractive contamination of $\mathcal{D}$ if a sample from $\mathcal{D}^{\prime}$ is equivalent to a sample from $\mathcal{D}$ conditioned on an event that occurs with probability at least $1-\eta$. We use $\mathrm{sub}_{\eta}(\mathcal{D})$ to denote the set of all such $\mathcal{D}^{\prime}$.
Similarly, for any $S\in X^{m}$, we say that $S^{\prime}$ is an $\eta$-subtractive contamination of $S$ if it is formed by removing at most $\lfloor\eta\cdot m\rfloor$ arbitrary points from $S$. In a slight overload of notation, we use $\mathrm{sub}_{\eta}(S)$ to denote the set of all such $S^{\prime}$.
We will show, as an easy consequence of Theorem 3, that the oblivious and adaptive variants of subtractive contamination are equivalent.
Theorem 10 (Oblivious and adaptive subtractive contamination are equivalent).
For any $\eta,\varepsilon\in(0,1)$, $f:X^{n}\to\{0,1\}$, and distribution $\mathcal{D}$ over $X$, let $M=\mathrm{poly}(n,1/\varepsilon,1/(1-\eta))$. Then,
$\left|\sup_{\mathcal{D}^{\prime}\in\mathrm{sub}_{\eta}(\mathcal{D})}\left(\mathop{{\mathds{E}}\/}_{\bm{S}\sim(\mathcal{D}^{\prime})^{n}}[f(\bm{S})]\right)-\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{M}}\left[\sup_{\bm{S}^{\prime}\in\mathrm{sub}_{\eta}(\bm{S})}\left(\mathop{{\mathds{E}}\/}[f\circ\Phi_{{\star}\to n}(\bm{S}^{\prime})]\right)\right]\right|\leq\varepsilon.$
Theorem 10 is an easy consequence of Theorem 6 as well as the below lemma.
Lemma A.1 (Converting the subtractive adversary to our framework).
For any $\eta,\varepsilon\in(0,1)$, $f:X^{n}\to\{0,1\}$, and distribution $\mathcal{D}$ over $X$, let
$m\coloneqq\left\lceil\frac{\max\left(2n,\sqrt{8\ln(1/\varepsilon)}\right)}{1-\eta}\right\rceil,$ (9)
and $X^{\prime}\coloneqq X\cup\{\varnothing\}$. There is a degree-$2$ cost function $\rho$ and $f^{\prime}:(X^{\prime})^{m}\to\{0,1\}$ for which
$\left|\sup_{\mathcal{D}^{\prime}\in\mathrm{sub}_{\eta}(\mathcal{D})}\left(\mathop{{\mathds{E}}\/}_{\bm{S}\sim(\mathcal{D}^{\prime})^{n}}[f(\bm{S})]\right)-\mathrm{Oblivious\text{-}Max}_{\rho^{\prime}}(f^{\prime},\mathcal{D})\right|\leq\varepsilon,$
and, for all $M\geq m$,
$\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{M}}\left[\sup_{\bm{S}^{\prime}\in\mathrm{sub}_{\eta}(\bm{S})}\left(\mathop{{\mathds{E}}\/}[f\circ\Phi_{{\star}\to n}(\bm{S}^{\prime})]\right)\right]=\mathrm{Adaptive\text{-}Max}_{\rho}(f^{\prime}\circ\Phi_{M\to m},\mathcal{D})$
Proof of Theorem 10 assuming Lemma A.1.
Let $f^{\prime}$, $\rho$, and $\mathcal{D}$ be as in Lemma A.1. By Theorem 6, for $M=\mathrm{poly}(m,1/\varepsilon)=\mathrm{poly}(n,1/(1-\eta),\varepsilon)$,
$\left|\mathrm{Oblivious\text{-}Max}_{\rho}(f^{\prime},\mathcal{D}^{\prime})-\mathrm{Adaptive\text{-}Max}_{\rho}(f^{\prime}\circ\Phi_{M\to m},\mathcal{D}^{\prime})\right|\leq\varepsilon.$
By the first part of Lemma A.1 and triangle inequality, we have that
$\left|\sup_{\mathcal{D}^{\prime}\in\mathrm{sub}_{\eta}(\mathcal{D})}\left(\mathop{{\mathds{E}}\/}_{\bm{S}\sim(\mathcal{D}^{\prime})^{n}}[f(\bm{S})]\right)-\mathrm{Adaptive\text{-}Max}_{\rho}(f^{\prime}\circ\Phi_{M\to m},\mathcal{D}^{\prime})\right|\leq 2\varepsilon.$
By the first part of Lemma A.1 this implies that
$\left|\sup_{\mathcal{D}^{\prime}\in\mathrm{sub}_{\eta}(\mathcal{D})}\left(\mathop{{\mathds{E}}\/}_{\bm{S}\sim(\mathcal{D}^{\prime})^{n}}[f(\bm{S})]\right)-\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{M}}\left[\sup_{\bm{S}^{\prime}\in\mathrm{sub}_{\eta}(\bm{S})}\left(\mathop{{\mathds{E}}\/}[f\circ\Phi_{{\star}\to n}(\bm{S}^{\prime})]\right)\right]\right|\leq 2\varepsilon.$
The desired result holds by renaming $\varepsilon^{\prime}=\varepsilon/2$. ∎
The proof of Lemma A.1 will use the following basic concentration inequality
Proposition A.2.
Let $\mathcal{D}$ be a distribution on $X^{\prime}\coloneqq X\cup\{\varnothing\}$ for which
$\mathop{{\operatorname{{Pr}}}\/}_{\bm{x}\sim\mathcal{D}^{\prime}}[\bm{x}=\varnothing]\leq\eta.$
For $m$ as in Equation 9,
$\mathop{{\operatorname{{Pr}}}\/}_{\bm{S}\sim(\mathcal{D}^{\prime})^{m}}\left[\sum_{x\in\bm{S}}\mathds{1}[x\neq\varnothing]\leq n\right]\leq\varepsilon.$
Proof.
Let $\bm{z}$ be the random variable that counts the number of entries $\bm{S}^{\prime}$ has that are not equal to $\emptyset$. Then,
$\mu\coloneqq\mathop{{\mathds{E}}\/}[\bm{z}]\geq\frac{m}{1-\eta}\geq\max\left(2n,\sqrt{8\ln(1/\varepsilon)}\right).$
Furthermore, $\bm{z}$ is the sum of independent random variables taking on values in $\{0,1\}$ (each indicating whether $\bm{S}_{i}\neq\emptyset$ for some index $i$). By a standard Chernoff bound (Fact 5.1),
$\mathop{{\operatorname{{Pr}}}\/}[\bm{z}\leq n]\leq e^{-\mu^{2}/8}\leq\varepsilon.$
Proof of Lemma A.1.
We begin by constructing the cost function. In the original subtractive adversary, for each $x$, the adversary can either choose to keep $x$ in the sample or remove it at a cost of $1/\eta$. We will construct $\rho$ so that the adversary has the same options and represent this “removal” option as converting an input $x$ to $\varnothing$:
$\rho(x,y)\coloneqq\begin{cases}0&\text{if }x=y\\ \frac{1}{\eta}&\text{if }x\neq y\text{ and }y=\varnothing\\ \infty&\text{otherwise}.\end{cases}$ (10)
The function $f^{\prime}$ simply runs $f$ on a random subset of its non-null input. For any $S\in X^{m}$, let $S_{\neq\varnothing}$ denote the subset of $S$ consisting of all points not equal to $\varnothing$. Then,
$f^{\prime}(S)\coloneqq\begin{cases}f(\Phi_{{\star}\to n}(S_{\neq\varnothing}))&\text{if $|S_{\neq\varnothing}|\geq n$}\\ 0&\text{otherwise.}\end{cases}$
Next, we analyze the oblivious adversaries. Consider a draw $\bm{x}\sim\mathcal{D}$ coupled to an event $\bm{E}$ occurring with probability at least $1-\delta$. Then, $\mathrm{sub}_{\eta}(\mathcal{D})$ consists of all possible distributions of $\bm{x}$ conditioned on $\bm{E}$, whereas, based on Equation 10, $\mathcal{C}_{\rho}(\mathcal{D})$ consists of all the possible distributions of $\bm{y}$ where
$\bm{y}\coloneqq\begin{cases}\bm{x}&\text{if $\bm{E}$}\\ \varnothing&\text{otherwise.}\end{cases}$
For any such event $\bm{E}$, let $\mathcal{D}_{1}^{\prime}\in\mathrm{sub}_{\eta}(\mathcal{D})$ and $\mathcal{D}_{2}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$ be the corresponding distribution. We will show that
$\left|\mathop{{\mathds{E}}\/}_{\bm{S}_{1}\sim(\mathcal{D}^{\prime}_{1})^{n}}[f(\bm{S}_{1})]-\mathop{{\mathds{E}}\/}_{\bm{S}_{2}\sim(\mathcal{D}^{\prime}_{2})^{m}}[f(\bm{S}_{2})]\right|\leq\varepsilon.$
Each element of $\bm{S}_{2}$ is set to $\varnothing$ with a probability that is at most $\eta$ and otherwise has the same distribution as an element of $\bm{S}_{1}$. Therefore, the above difference is bounded by the probability that $\bm{S}_{2}$ less than $n$ non-null elements. This is at most $\varepsilon$ by Proposition A.2.
For the adaptive equivalence, consider any $S\in X^{M}$. Then any $S_{1}\in\mathrm{sub}_{\eta}(S)$ is formed by removing at most $\lfloor\eta M\rfloor$ of the points in $S$, whereas $S_{2}\in\mathcal{C}_{\rho}(S)$ is formed by setting $\lfloor\eta M\rfloor$ of the points to $\varnothing$. Suppose we remove the same set of points to form $S_{1}$ as we set to $\varnothing$ to form $S_{2}$. Then, using the fact that at least $n$ points must remain unchanged since $M\geq m\geq n/(1-\eta)$ and that the subsampling filter composes
$\mathop{{\mathds{E}}\/}[f\circ\Phi_{{\star}\to n}(S_{1})]=\mathop{{\mathds{E}}\/}[f^{\prime}\circ\Phi_{M\to m}(S_{2})].$
Hence, for any choice of $S\in X^{M}$,
$\sup_{S_{1}\in\mathrm{sub}_{\eta}(S)}\left(\mathop{{\mathds{E}}\/}[f\circ\Phi_{{\star}\to n}(S_{1})]\right)=\sup_{S_{2}\in\mathcal{C}_{\rho}(S)}\left(\mathop{{\mathds{E}}\/}[f^{\prime}\circ\Phi_{M\to m}(S_{2})]\right).$
This implies the desired result. ∎
A.2 Additive adversaries
First, we formally define $\eta$-additive corruptions. Note that the oblivious adversary below exactly corresponds to Huber’s original contamination model [Hub64].
Definition 16 (Standard additive adversaries).
For any distribution $\mathcal{D}$ and $\eta\in(0,1)$, we say that $\mathcal{D}^{\prime}$ is an $\eta$-additive contamination of $\mathcal{D}$ if, for some distribution $\mathcal{E}$,
$\mathcal{D}^{\prime}\coloneqq(1-\eta)\mathcal{D}+\eta\mathcal{E}$
We use $\mathrm{add}_{\eta}(\mathcal{D})$ to denote the set of all such $\mathcal{D}^{\prime}$, and for any function $f:X^{n}\to\{0,1\}$, define
$\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})\coloneqq\sup_{\mathcal{D}^{\prime}\in\mathrm{add}_{\eta}(\mathcal{D})}\left\{\mathop{{\mathds{E}}\/}_{\bm{S}^{\prime}\sim(\mathcal{D}^{\prime})^{n}}(f(\bm{S}^{\prime})\right\}.$
Similarly, for any $S\in X^{\lceil(1-\eta)m\rceil}$, we say $S^{\prime}$ is an $\eta$-additive contamination of $S$ if it is formed by adding $\lfloor\eta m\rfloor$ points to $S$ and then arbitrarily permuting it. In a slight overload of notation, we use $\mathrm{add}_{\eta}(S)$ to denote the set of all such $S^{\prime}$, and for any $f:X^{m}\to\{0,1\}$, write
$\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})\coloneqq\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{\lceil(1-\eta)m\rceil}}\left[\sup_{\bm{S}^{\prime}\in\mathrm{add}_{\eta}(\bm{S})}\left(\mathop{{\mathds{E}}\/}[f(\bm{S}^{\prime})]\right)\right].$
The equivalence between these two adversaries was already shown by [BLMT22], but we also prove it here to show our framework can recover their result.
Theorem 11 (Oblivious and adaptive additive contamination are equivalent).
For any $\eta,\varepsilon\in(0,1)$, $f:X^{n}\to\{0,1\}$, and distribution $\mathcal{D}$ over $X$, let $m=\mathrm{poly}(n,1/\varepsilon,\ln|X|)$. Then,
$\left|\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})-\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f\circ\Phi_{{m}\to n},\mathcal{D})\right|\leq\varepsilon.$
To prove this equivalence, we will introduce a variant of the adaptive adversary, the binomial adversary. In this variant, rather than drawing exactly $\lceil(1-\eta)m\rceil$ clean points and the adversary being able to add $\lfloor\eta m\rfloor$ corrupted points, the number of clean points is itself drawn randomly from a binomial distribution.
Definition 17 (Binomial adversary).
For any distribution $\mathcal{D}$ and sample size $n$, the binomial adversary first draws $\bm{z}\sim\mathrm{Bin}(n,(1-\eta))$ clean points from $\mathcal{D}$, then adds $n-\bm{z}$ arbitrary points to this clean sample, and finally permutes all $m$ points arbitrarily. For any $f:X\to\{0,1\}$, we define
$\mathrm{Binomial\text{-}Max}_{\eta}(f,\mathcal{D})\coloneqq\mathop{{\mathds{E}}\/}_{\bm{z}\sim\mathrm{Bin}(n,1-\eta)}\left[\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{\bm{z}}}\left[\sup_{\bm{S}^{\prime}\in\mathrm{complete}_{m}(\bm{S})}\left(\mathop{{\mathds{E}}\/}[f(\bm{S}^{\prime})]\right)\right]\right]$
where $\mathrm{complete}_{n}(S)$ to denote all samples that can be created by adding $n-|S|$ points to $S$.
Our proof of Theorem 11 proceeds in two steps. We first use Theorem 2 to show that the oblivious additive adversary is equivalent to the binomial adversary, and then show equivalence between the binomial adversary and adaptive additive adversary.
Proposition A.3 (The oblivious adversary and binomial adversary are equivalent).
For any $\eta,\varepsilon\in(0,1)$, $f:X^{n}\to\{0,1\}$, and distribution $\mathcal{D}$ over $X$, let $m=\mathrm{poly}(n,1/\varepsilon,\ln|X|)$. Then,
$\left|\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})-\mathrm{Binomial\text{-}Max}_{\eta}(f\circ\Phi_{{m}\to n},\mathcal{D})\right|\leq\varepsilon.$
Proof.
This will be a fairly straightforward application of Theorem 2. Let $X^{\prime}\coloneqq X\cup\{\varnothing\}$ and $\mathcal{D}_{\eta}$ be the distribution that outputs $\varnothing$ with probability $\eta$ and otherwise outputs $\mathcal{D}$,
$\mathcal{D}_{\eta}\coloneqq(1-\eta)\cdot\mathcal{D}+\eta\cdot\{\varnothing\}.$
Then, we’ll define an adversary that can send $\varnothing$ to any element of $X^{\prime}$ but otherwise cannot change its input.
$\rho(x,y)\coloneqq\begin{cases}0&\text{if $x=y$ or $y=\varnothing$}\\ \infty&\text{otherwise.}\end{cases}$
Finally, let $f^{\prime}:X^{n}\to\{0,1\}$ be defined as
$f^{\prime}(S)\coloneqq\begin{cases}0&\text{if $\varnothing\in S$}\\ f(S)&\text{otherwise.}\end{cases}$
We observe that,
$\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})$ $=\mathrm{Oblivious\text{-}Max}_{\rho}(f^{\prime},\mathcal{D}_{\eta}),\quad\text{and}$
$\mathrm{Binomial\text{-}Max}_{\eta}(f\circ\Phi_{{m}\to n},\mathcal{D})$ $=\mathrm{Adaptive\text{-}Max}_{\rho}(f^{\prime}\circ\Phi_{m\to n},\mathcal{D}_{\eta}).$
because in order to maximize the success probability of $f^{\prime}$, the adversaries should send every $\varnothing$ they see to their adversarial choice of an element in $X$. The desired result follows from Theorem 6. ∎
Next, we show that the binomial adversary and adaptive additive adversary are equivalent.
Proposition A.4 (The binomial and adaptive additive adversaries are equivalent).
For any $f:X^{n}\to\{0,1\}$, $\varepsilon,\eta\in(0,1)$, and distribution $\mathcal{D}$, let
$m=O\left(\frac{n^{2}}{\varepsilon^{2}}\right).$
Then, for $f^{\prime}=f\circ\Phi_{{m}\to n}$
$\left|\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})-\mathrm{Binomial\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})\right|\leq\varepsilon.$
Proof.
Expanding the definitions, we wish to show that,
$\left|\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{\lceil(1-\eta)m\rceil}}\left[\sup_{\bm{S}_{1}^{\prime}\in\mathrm{add}_{\eta}(\bm{S})}\left(\mathop{{\mathds{E}}\/}[f^{\prime}(\bm{S}_{1}^{\prime})]\right)\right]-\mathop{{\mathds{E}}\/}_{\bm{z}\sim\mathrm{Bin}(m,1-\eta)}\left[\mathop{{\mathds{E}}\/}_{\bm{S}\sim\mathcal{D}^{\bm{z}}}\left[\sup_{\bm{S}_{2}^{\prime}\in\mathrm{complete}_{m}(\bm{S})}\left(\mathop{{\mathds{E}}\/}[f^{\prime}(\bm{S}_{2}^{\prime})]\right)\right]\right]\right|\leq\varepsilon.$
Regardless of the strategy of one adversary, it is possible to choose a strategy for the other adversary so that the following holds: There is a coupling of $\bm{S}_{1}^{\prime}$ and $\bm{S}_{2}^{\prime}$ for which the expected number of differences between $\bm{S}_{1}^{\prime}$ and $\bm{S}_{2}^{\prime}$ is at most $\mathop{{\mathds{E}}\/}\left[\left|\bm{z}-\lceil(1-\eta)m\rceil\right|\right]$. Furthermore, for any $S_{1},S_{2}$ differing in at most $\Delta$ points and function $f:X^{n}\to\{0,1\}$,
$\left|\mathop{{\mathds{E}}\/}[f\circ\Phi_{{m}\to n}(S_{1})]-\mathop{{\mathds{E}}\/}[f\circ\Phi_{{m}\to n}(S_{2})]\right|\leq\frac{n\Delta}{m}=\varepsilon/2,$
because, in order for the two above quantities to differ, the subsample must select one of the $\Delta$ differences. Therefore,
$\left|\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})-\mathrm{Binomial\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})\right|\leq\frac{n}{m}\cdot\mathop{{\mathds{E}}\/}\left[\left|\bm{z}-\lceil(1-\eta)m\rceil\right|\right]\leq O\left(\frac{n}{\sqrt{m}}\right)\leq\varepsilon.\qed$
Finally, we prove the main result of this section.
Proof of Theorem 11.
Let $f^{\prime}=f\circ\Phi_{{m}\to n}$. Then, by triangle inequality
$\left|\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})-\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})\right|$
$\quad\quad\leq\left|\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})-\mathrm{Binomial\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})\right|$
$\quad\quad+\left|\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})-\mathrm{Binomial\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})\right|.$
Each of the above terms is at most $\varepsilon/2$ by Propositions A.3 and A.4. ∎
Appendix B Partially-adaptive adversaries
In this section, we introduce two partially adaptive adversaries, malicious noise and the non-iid adversary, and prove that both are equivalent to fully adaptive and fully oblivious adversaries.
Definition 18 (Malicious noise, [Val85]).
In malicious noise with base distribution $\mathcal{D}$ and noise rate $\eta$, the sample $\bm{S}=(\bm{x}_{1},\ldots,\bm{x}_{n})$ is generated sequentially. For each $i\in[n]$, an $\eta$-coin is flipped and then,
  1. 1.
    If the coin is tails, a clean point is sampled $\bm{x}_{i}\sim\mathcal{D}$.
  2. 2.
    If the coin is heads, the adversary gets to choose $\bm{x}_{i}$ arbitrarily with full knowledge of $\bm{x}_{1},\ldots,\bm{x}_{i-1}$ but no knowledge of the future points ($\bm{x}_{i+1},\ldots,\bm{x}_{n})$.
For any $f:X^{n}\to\{0,1\}$ and distribution $\mathcal{D}$, we’ll use $\mathrm{Mal\text{-}Max}_{\eta}(f,\mathcal{D})$ to denote the maximum expected value of $f(\bm{S})$ over any $\bm{S}$ generated by a malicious adversary with noise rate $\eta$.
The malicious adversary is partially adaptive in the sense that, when it chooses how to corrupt $\bm{x}_{i}$, it only knows the points generated in the past.
Definition 19 (The non-iid adversary, [CHL+23]).
In the non-iid adversary model with base distribution $\mathcal{D}$ and noise rate $\eta$, to generate $n$ samples, first the adversary arbitrarily chooses $\lfloor\eta n\rfloor$ points, and then $\lceil(1-\eta)n\rceil$ points are generated iid from $\mathcal{D}$, added to the generated points, and permuted arbitrarily. For any $f:X^{n}\to\{0,1\}$ and distribution $\mathcal{D}$, we’ll use $\mathrm{Non\text{-}iid\text{-}Max}_{\eta}(f,\mathcal{D})$ to denote the maximum expected value of $f(\bm{S})$ over any $\bm{S}$ generated by a non-iid adversary with noise rate $\eta$.
This adversary is referred to as non-iid because the $\lfloor\eta n\rfloor$ points can be generated arbitrarily and need not be iid. If they were, this adversary would be extremely similar to the oblivious additive adversary, with the only difference being that the non-iid adversary generates exactly $\lfloor\eta n\rfloor$ corruptions, whereas the oblivious adversary generates $\mathrm{Bin}(n,\eta)$ corruptions.
Theorem 12 (Equivalence of all additive adversaries).
For any $\eta,\varepsilon\in(0,1)$, $f:X^{n}\to\{0,1\}$, and distribution $\mathcal{D}$ over $X$, let $m=\mathrm{poly}(n,1/\varepsilon,\ln|X|)$ and $f^{\prime}=f\circ\Phi_{{m}\to n}$. The following are all within ±$\pm\varepsilon$ of one another.
  1. 1.
    The maximum success probability of the oblivious additive adversary,
    $\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D}).$
  2. 2.
    The maximum success probability of the adaptive additive adversary,
    $\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f^{\prime},\mathcal{D}).$
  3. 3.
    The maximum success probability of the malicious adversary,
    $\mathrm{Mal\text{-}Max}_{\eta}(f^{\prime},\mathcal{D}).$
  4. 4.
    The maximum success probability of the non-iid adversary,
    $\mathrm{Non\text{-}iid\text{-}Max}_{\eta}(f^{\prime},\mathcal{D}).$
We already proved that $\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})$ and $\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})$ are within ±$\pm\varepsilon$ of one another. To prove the same for malicious noise, we will show that malicious noise is no more powerful than the adaptive adversary, and at least as powerful as the oblivious adversary.
Proposition B.1.
In the setting Theorem 12,
$\mathrm{Mal\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})\leq\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})+\varepsilon.$
Proof.
Consider an arbitrary malicious adversary. This adversary can make $\bm{z}$ many corruptions, where $\bm{z}\sim\mathrm{Bin}(m,(1-\eta))$. Therefore, if the malicious adversary knew the full sample, it would be the same adversary as the binomial adversary. As a result, for any choices of the malicious adversary, there is a binomial adversary simulating it (generating the same distribution over corrupted samples). This gives that
$\mathrm{Mal\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})\leq\mathrm{Binomial\text{-}Max}_{\eta}(f^{\prime},\mathcal{D}).$
The desired result then follows from Proposition A.4. ∎
Proposition B.2.
In the setting of Theorem 12,
$\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})\leq\mathrm{Mal\text{-}Max}_{\eta}(f^{\prime},\mathcal{D}).$
Proof.
Consider any strategy for the oblivious adversary. It chooses an arbitrary distribution $\mathcal{E}$ and sets
$\mathcal{D}^{\prime}=(1-\eta)\cdot\mathcal{D}+\eta\cdot\mathcal{E}.$
Now, consider the malicious adversary that, whenever it can corrupt a point, it draws a point from $\mathcal{E}$ as its corruption. Then, each of the $m$ points in this malicious adversary’s sample is independent and drawn from $\mathcal{D}^{\prime}$. After subsampling uniformly without replacement, the $n$ points will still be independent and drawn from $\mathcal{E}$. This means that for any choices of the oblivious adversary, the malicious adversary can simulate them, giving the desired inequality. ∎
We execute the same two steps for the non-iid adversary.
Proposition B.3.
In the setting Theorem 12,
$\mathrm{Non\text{-}iid\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})\leq\mathrm{Adaptive\text{-}Add\text{-}Max}_{\eta}(f^{\prime},\mathcal{D}).$
Proof.
Consider any strategy for the non-iid adversary. This is a set of points $x_{1},\ldots,x_{\lfloor\eta m\rfloor}$ it will add to the sample. Now, consider the adaptive adversary that adds these same points regardless of what the clean points are. It’s straightforward to see this adaptive adversary simulates the non-iid adversary, giving the desired inequality. ∎
Proposition B.4.
In the setting Theorem 12,
$\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})\leq\mathrm{Non\text{-}iid\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})+\varepsilon.$
Proof.
Consider any strategy for the oblivious adversary. It chooses an arbitrary distribution $\mathcal{E}$ and sets
$\mathcal{D}^{\prime}=(1-\eta)\cdot\mathcal{D}+\eta\cdot\mathcal{E}.$
To draw a sample $\bm{S}\sim(\mathcal{D}^{\prime})^{n}$, we can first independently draw indicators $\bm{a}_{1},\ldots,\bm{a}_{n}\overset{\mathrm{iid}}{\sim}\mathrm{Ber}(\eta)$. For every $i$ in which $\bm{a}_{i}=1$, $\bm{S}_{i}$ is sampled from $\mathcal{E}$. In contrast, if $\bm{a}_{i}=0$, then $\bm{S}_{i}$ is sampled from $\mathcal{D}$.
Now consider the non-iid adversary that draws $\bm{x}_{1},\ldots,\bm{x}_{\lfloor\eta m\rfloor}\overset{\mathrm{iid}}{\sim}\mathcal{E}$ and adds them to the sample, and let ${\bm{T}}$ be a size-$n$ subsample from the resulting non-iid adversaries subsample. Let $\bm{b}_{i}$ be the indicator for whether the $i^{\text{th}}$ element of ${\bm{T}}$ comes from one of these $\lfloor\eta m\rfloor$ points added. Then, we observe that, after conditioning on the values of $\bm{b}_{1},\ldots,\bm{b}_{n}$, each element of ${\bm{T}}$ is independently drawn from $\mathcal{E}$ if $\bm{b}_{i}=1$ and $\mathcal{D}$ otherwise. We observe that the distribution of $(\bm{b}_{1},\ldots,\bm{b}_{n})$ is equivalent to the distribution obtained by first drawing $\bm{i}_{1},\ldots,\bm{i}_{n}$ uniformly without replacement from $[m]$ and then setting $\bm{b}_{i}=\mathds{1}[\bm{i}_{i}\leq\lfloor\eta n\rfloor]$.
Therefore, the desired result follows from showing that the TV distance of the distributions of $\bm{a}$ and $\bm{b}$ is at most $\varepsilon$. We prove by exhibiting a coupling of $\bm{a}$ and $\bm{b}$ for which they differ with probability at most $\varepsilon$.
  1. 1.
    Draw $\bm{z}_{1},\ldots\bm{z}_{n}$ uniformly and independently $[0,1]$.
  2. 2.
    Set $\bm{a}_{j}=\mathds{1}[\bm{z}_{j}\leq\eta]$ for each $j\in[n]$.
  3. 3.
    For each $j\in[n]$, let $\bm{i}_{n}=\lfloor\bm{z}_{i}\cdot m\rfloor+1$. Note this gives that $\bm{i}_{1},\ldots,\bm{i}_{n}$ are each uniform on $[m]$ and they are independent.
  4. 4.
    If $\bm{i}_{1},\ldots,\bm{i}_{n}$ are not unique, resample them by drawing them uniformly from $[m]$ without replacement.
  5. 5.
    Set $\bm{b}_{j}=\mathds{1}[\bm{i}_{j}\leq\lfloor\eta m\rfloor]$.
First, we confirm that the marginal distributions are correct. Each $\bm{a}_{j}$ is independent and drawn from $\mathrm{Ber}(\eta)$, so $\bm{a}$ has the correct marginal distribution.
If we resample, then $\bm{i}_{1},\ldots,\bm{i}_{n}$ are a uniform set of $n$ distinct indices from $[m]$. If we don’t resample, then they are also a uniform set of $n$ distinct indices from $[m]$, because before resampling, they are independent and uniform from $[m]$, and we only don’t resample if they are distinct. Therefore, the marginal distribution of $\bm{b}$ is correct.
Finally, we bound the probability $\bm{a}\neq\bm{b}$. There are two ways that $\bm{a}$ and $\bm{b}$ could be different.
  1. 1.
    One the $\bm{z}_{i}$ is between $\frac{\lfloor\eta m\rfloor}{m}$ and $\eta$. This occurs with probability at most $\frac{1}{m}$.
  2. 2.
    We needed to resample $\bm{i}_{1},\ldots,\bm{i}_{n}$ because they were not unique. This occurs if $\bm{i}_{j}=\bm{i}_{k}$ for $j\neq k$. By union bound, it occurs with probability at most $\frac{\binom{n}{2}}{m}\leq n^{2}/m$.
Union bounding over the above two, we have that
$\mathrm{Oblivious\text{-}Add\text{-}Max}_{\eta}(f,\mathcal{D})\leq\mathrm{Non\text{-}iid\text{-}Max}_{\eta}(f^{\prime},\mathcal{D})+\frac{1}{m}+\frac{n^{2}}{m}.\qed$
Theorem 12 is immediate from Propositions B.1, B.2, B.3 and B.4 and Theorem 11.
Appendix C Brief overview of [BLMT22]’s approaches and their limitations
C.1 The special case of additive adversaries
[BLMT22] proved that oblivious and adaptive additive adversaries are equivalent (corresponding to Theorem 11). Here we briefly describe their proof strategy, and why it does not generalize to other adversary models. For simplicity, we set $\eta=1/2$ in the below exposition.
Recall that the adaptive additive adversary, given a sample $S\in X^{M/2}$ can construct $S\cup T$ for arbitrary $T\in X^{M/2}$. Using a standard concentration inequality, for any $f:X^{n}\to\{0,1\}$ fixed choice of the corruption $T$,
$\mathop{{\operatorname{{Pr}}}\/}_{\bm{S}\sim\mathcal{D}^{{}^{(1-\eta)\cdot M}}}\left[f\circ\Phi_{{M}\to n}(\bm{S}\cup T)>\mathrm{Oblivious\text{-}Max}+\varepsilon\right]\leq 2^{-\Omega_{n,\varepsilon}(M)}$ (11)
where $\mathrm{Oblivious\text{-}Max}$ is appropriately defined for the setting. At first glance, the number of choices for $T$ is $|X|^{M/2}$, which also grows exponentially in $M$. This makes it impossible to union bound over all choices of $T$. [BLMT22]’s key observation is that the space of possible corruptions (choices of $T$) can be easily discretized to one that is much smaller.
In particular, given any $T\in X^{M/2}$, let ${\bm{T}}$ be formed by,
  1. 1.
    First taking $m\leq M$ samples $\bm{x}_{1},\ldots,\bm{x}_{m}\overset{\mathrm{iid}}{\sim}\mathrm{Unif}(T)$.
  2. 2.
    Then, for $k\coloneqq\frac{M}{2m}$, constructing ${\bm{T}}$ by taking $k$ copies of each of $\bm{x}_{1},\ldots,\bm{x}_{m}$.
Then, $\Phi_{{M}\to n}(T)$ and $\Phi_{{M}\to n}({\bm{T}})$ look identical unless a collusion occurs (i.e. the same $\bm{x}_{i}$ is sampled twice). If $m=n^{2}/\varepsilon$, that collusion occurs with probability at most $\varepsilon$, which is negligible. Furthermore, there are only $|X|^{m}$ possible choices for ${\bm{T}}$, which does not grow exponentially with $M$. Therefore, the desired result can proven using Equation 11 and an appropriate concentration inequality.
This strategy crucially relies on the fact that, while the adaptive additive adversary can choose its corruption (choice of $T$) as a function of the clean sample $S$, the set of possible corruptions does not depend on $S$. Hence, adaptivity is inherently weaker for the additive adversary than for other models where the space of possible corruptions depends on the clean sample.
For example, consider the case of subtractive adversaries, where the adaptive adversary can remove half the points in the sample. For a sample $S\in X^{M}$, there are $\approx 2^{M}$ ways to remove half the points, each parameterized by a bit string $b\in\{0,1\}^{M}$ where $b_{i}$ indicates whether the $i^{\text{th}}$ point is removed. Crucially, the “effect” of a bit string $b$ depends on the clean sample $S$ — in the sense that for the adversary to determine whether removing $S_{i}$ is a good idea, it must know the value of $S_{i}$. In particular, if the adversary is only allowed to choose $b$ from a subset $B\subseteq\{0,1\}^{M}$ of size much smaller than $2^{M}$ that is fixed before seeing the clean sample $S$, the power of the adversary is greatly diminished. This makes it not clear how a similar discretization argument as [BLMT22] used for additive adversaries would work.
C.2 The special case of statistical query algorithms
[BLMT22] also approved the equivalence between oblivious and adaptive adversaries for algorithms that never directly examine their dataset and only access it through statistical queries (SQ) [Kea98].
Basics of the SQ framework. A SQ is a pair $(\varphi,\tau)$ where $\varphi:X\to[0,1]$ is the query and $\tau>0$ is the tolerance. For any distribution $\mathcal{D}$, a valid response to the query $(\varphi,\tau)$ is any value that is within ±$\pm\tau$ of $\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}[\varphi(\bm{x})]$. An SQ algorithm $A$ using $k$ queries of tolerance $\tau$ specifies a sequence of $k$ adaptively chosen queries, $(\varphi_{1},\tau),\ldots,(\varphi_{k},\tau)$. For each $t\in[k]$, it receives a response $v_{t}$ which is within ±$\pm\tau$ of $\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}[\varphi_{t}(\bm{x})]$, and the identity of $\varphi_{t+1}$ is allowed to depend on the prior response $v_{1},\ldots,v_{t}$. After receiving all responses $v_{1},\ldots,v_{k}$, the $A$ chooses an output $y\in Y$.
We say $y\in Y$ is a valid output of $A$ on distribution $\mathcal{D}$ if it is a response that $A$ can generate given valid responses $v_{1},\ldots,v_{k}$ which are each within ±$\pm\tau$ of $\mathop{{\mathds{E}}\/}_{\bm{x}\sim\mathcal{D}}[\varphi_{t}(\bm{x})]$. We can now state [BLMT22]’s main result for SQ algorithms.
Fact C.1 (Oblivious and adaptive adversaries are equivalent for the SQ framework).
Let $A$ be any SQ algorithm making $k$ queries of tolerance $\tau$, and $\rho$ be any cost function. For $m=\mathrm{poly}(k,\tau)$, there is an algorithm $A^{\prime}:X^{m}\to Y$ with the following guarantee: For any distribution $\mathcal{D}$ over $X^{m}$, draw $\bm{S}\sim\mathcal{D}^{m}$ and let an adversary choose $\bm{S}^{\prime}\in\mathcal{C}_{\rho}(\bm{S})$. Then, $A^{\prime}(\bm{S}^{\prime})$ is a valid output of $A$ on distribution $\mathcal{D}^{\prime}$ for some $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$ with high probability over the randomness of $\bm{S}$.
To understand the utility of Fact C.1, suppose we have an SQ algorithm $A$ that solves some task in the presence of an oblivious adversary. This means that, for all $\mathcal{D}^{\prime}\in\mathcal{C}_{\rho}(\mathcal{D})$, any valid output of $A$ on $\mathcal{D}^{\prime}$ is a good answer for this task. Then, Fact C.1 says that $A^{\prime}$ given an adaptively corrupted sample will also provide a good answer for this task with high probability.
One straightforward weakness of this result compared to ours is not every task that admits an efficient solution also admits an efficient solution by an SQ algorithm [BFJ+94]. Even for tasks that can be cast into the SQ framework, our result has advantages.
  1. 1.
    The SQ equivalence in Fact C.1 is not black box. Given an algorithm $A$ not already in the SQ framework, in order to design an $A^{\prime}$ that defeats adaptive adversaries, first the algorithm designer must find an SQ algorithm that is “equivalent” to $A$ in order to apply Fact C.1, a task that is not always trivial. In contrast, our result gives a black-box technique, via subsampling, to construct $A^{\prime}$.
  2. 2.
    The SQ equivalence in Fact C.1 does not have a well-defined sample overhead. Even if an algorithm $A:X^{n}\to Y$ can be cast into some $A_{\mathrm{SQ}}$ operating in the SQ framework, the number of queries and tolerance $A_{\mathrm{SQ}}$ needs is not a predicable function of $n$. Therefore, it’s unclear how much larger the $m$ in Fact C.1 will be than $n$. In contrast, Theorem 3 gives a simple expression for what $m$ is needed as a function of $n$.